### v2

# Critical Nodes Identification in Complex Networks: A Survey

Abstract
Complex networks have become essential tools for understanding diverse phenomena in social systems, traffic systems, biomolecular systems, and financial systems. Identifying critical nodes is a central theme in contemporary research, serving as a vital bridge between theoretical foundations and practical applications. Nevertheless, the intrinsic complexity and structural heterogeneity characterizing real-world networks, with particular emphasis on dynamic and higher-order networks, present substantial obstacles to the development of universal frameworks for critical node identification. This paper provides a comprehensive review of critical node identification techniques, categorizing them into seven main classes: centrality, critical nodes deletion problem, influence maximization, network control, artificial intelligence, higher-order and dynamic methods. Our review bridges the gaps in existing surveys by systematically classifying methods based on their methodological foundations and practical implications, and by highlighting their strengths, limitations, and applicability across different network types. Our work enhances the understanding of critical node research by identifying key challenges, such as algorithmic universality, real-time evaluation in dynamic networks, analysis of higher-order structures, and computational efficiency in large-scale networks. The structured synthesis consolidates current progress and highlights open questions, particularly in modeling temporal dynamics, advancing efficient algorithms, integrating machine learning approaches, and developing scalable and interpretable metrics for complex systems.
Keywords:
Complex Network, Critical Nodes, Network Structure, Complex System
Figure 1: Graphic Abstract.
1 Introduction
Complex networks and systems are fundamental to diverse fields, including social network analysis, urban traffic modeling communicability2018 , biomolecular structure analysis de2023more , and financial system bardoscia2021physics , where understanding the interplay between network structure, system architecture, and functionality is a central research focus Yu2024 . A typical network consists of nodes, representing entities within a real system, and edges, which encode relationships between them. By leveraging graph structures to model real-world systems, researchers can systematically analyze the interactions that govern their dynamics. A key challenge in complex network research is identifying important nodes, a topic that has garnered significant attention due to its broad applications. In neural brain networks, neurons are interconnected through nerve fibers, and the transmission of signals enables autonomous decision-making. In transportation networks, identifying critical nodes aids in traffic prediction, optimization, and infrastructure planning. Similarly, in epidemiological networks, infections originating from a few individuals can rapidly propagate through contact links, posing significant risks to public health. In communication networks, detecting key nodes enhances resource allocation and strengthens network security. Likewise, in social networks, pinpointing influential individuals or communities facilitates targeted marketing and improves user engagement. Advances in structural analysis and feature extraction continue to refine our understanding of network robustness, efficiency, and optimization, shedding light on the fundamental principles that govern complex systems.
Table 1: Comparison of existing studies on critical node identification across different methodological aspects.
Centrality CNDP IM Control AI Higher-order Dynamic
Liu et al. REN2013
Ren et al. xiaolong2014
Lalou et al. lalou2018critical
Hafiene et al. hafiene2020influential
Zhang et al. Zhang2023EpidemicMN
Li et al. li2023survey
Jaouadi et al. jaouadi2024survey
Liu et al. Liu2024Fundamental
Chen et al.(Ours)
Recent years have seen a growing number of review articles providing systematic overviews of node importance in complex networks. Various terms—such as key nodes, critical nodes, important nodes, and influence-maximizing nodes—have been used to describe structurally and functionally significant elements, leading to the development of a wide range of methods for assessing node importance in different contexts. For instance, Liu et al. REN2013 reviewed advances in node influence ranking by considering both network structure and diffusion dynamics, analyzing the strengths, limitations, and applicability of different approaches. Ren et al. xiaolong2014 introduced over 30 methods for identifying important nodes, categorizing them based on criteria such as node neighbors, network paths, feature vectors, and contraction techniques. Their study evaluated node importance from the perspectives of network robustness, vulnerability, and diffusion dynamics, yet the effectiveness of these methods remains constrained by network size and computational complexity. Lalou et al. lalou2018critical focused on the Critical Node Deletion Problem (CNDP), classifying solution algorithms based on problem variants and constraints related to network fragmentation. While their work systematically examined the impact of node removal on network connectivity and vulnerability, its applicability to dynamic and real-world networks remains an open challenge. More recent reviews have explored influence maximization (IM) and node ranking in greater depth. Hafiene et al. hafiene2020influential addressed the IM problem in social networks, aiming to select a set of users that maximize influence spread. However, existing IM approaches often struggle with high computational costs and scalability in large-scale networks. Zhang et al. Zhang2023EpidemicMN classified influential node ranking methods (INRM) and compared validation models based on cascade failure, linear threshold, and epidemic processes, yet integrating these models with real-time applications remains a challenge. Li et al. li2023survey summarized key challenges in diffusion models, such as computational complexity and parameter tuning, and examined how reinforcement learning and graph learning improve efficiency and generalizability. Nevertheless, machine learning–based approaches often require extensive training data and careful model selection. Jaouadi et al.  jaouadi2024survey provided a comprehensive classification of IM models, highlighting their evolution from static to dynamic networks and emerging trends in handling large-scale systems, though trade-offs between accuracy and efficiency remain a concern. In addition to traditional network models, recent efforts have extended node importance analysis to high-order networks. Liu et al. Liu2024Fundamental reviewed statistical indices for hypergraphs and simplicial complexes, introducing physically meaningful measures such as degree-dependent indicators, clustering coefficients, centrality metrics, and entropy-based indices. These studies highlight the diversity of methods for evaluating node importance, as well as significant progress in influence propagation, network robustness, and higher-order structures. However, most previous reviews have focused on specific types of network structures, such as static or higher-order networks, and there is still a lack of comprehensive and unified analyses covering multiple network forms. From the perspective of application scenarios, including influence maximization and network fragmentation, the importance of nodes in relation to network control and stability has been largely overlooked. Similarly, from a methodological perspective, while centrality measures, node deletion strategies, and infection models have been discussed, systematic reviews of machine learning-based approaches and comprehensive index-based methods remain insufficient. This paper provides a comprehensive review of recent advances in identifying critical nodes in complex networks. Our key contributions are as follows:
Figure 2: Taxonomy of Critical Node Identification Methods in Complex Networks.
Table 2: Overview of Structure, Categorization and Contributions.
Ranking Methods Categories Details and Contributions
Neighbors-Based Degree Centrality Focus on local structural features, including node degrees, coreness, neighbor influence, and cohesive subgraphs, providing intuitive and efficient influence estimation.
K-shell Decomposition
H-index
K-truss
Structural Hole
Neighborhood Similarity
Eigenvector-Based Eigenvector Centrality Utilize spectral properties and iterative propagation principles to assess global influence, emphasizing connections to other influential nodes.
Katz Centrality
Pagerank Algorithm
Voting Mechanism
Path-Based Betweenness Centrality Measure node importance based on shortest paths, traversal distances, or probabilistic walks, capturing both local and global topological information.
Closeness Centrality
Random Walk
Control-Optimization Network Control Formulate node selection as optimization or control problems, identifying key nodes for structural stability, synchronization, or influence maximization.
Maximization of Connected Component
Minimization of Pair Connectivity
Minimum Spanning Tree
Machine Learning-Based Information Entropy Leverage machine learning models, including classical, deep, and reinforcement learning, to learn influence patterns from structural or dynamic features.
Clustering Coefficient
Machine Learning
Reinforcement Learning
Comprehensive Index Gravity Formula Integrate multiple metrics or criteria into unified indices, achieving more balanced and adaptable node ranking by considering multiple dimensions of importance.
Multiple Metrics
Topsis and Entropy-Weight
Information Propagation Diffusion Model Estimate node influence through simulation or modeling of propagation dynamics, linking rankings to spreading.
Dynamical Sensitivity
High-order Networks Capture complex relational patterns beyond pairwise links to improve influence identification accuracy .
Dynamic Networks Address temporal evolution of networks to track time-dependent variations in node importance .
  • Categorization of Application Scenarios: From the application scenario oriented perspective of key nodes, this study systematically classifies existing methods into three key categories as depicted in Figure. 2: (1) topological representation, which assesses node importance by structural metrics such as degree, eigenvalues, and path-based centrality; (2) influence maximization, where the goal is to select a set of Top-$k$ influential seed nodes, $\{v_{1},v_{2},\dots,v_{k}\}$, to maximize information spread under diffusion models; and (3) network robustness examines the impact of node removal on network connectivity, modeled as an optimization problem with an objective function $f(G\setminus S)$, where $S\subseteq V$ represents the set of removed nodes.
  • Bridging Gaps in Existing Surveys: Despite prior studies focusing on specific aspects such as centrality, influence propagation, and critical node deletion problems, there exists a lack of unified discourse regarding the methods and application scenarios of machine learning techniques in recent years. As presented in Table. 1, this study bridges this gap by integrating artificial intelligence methods. These methods will be classified based on their methodological foundations and practical applications. Via comparative analysis, the study underscores the advantages, limitations, and applicability of these methods across different network types.
  • Comprehensive Review of Methodologies: From a methodological standpoint, this study analyzes nine key node identification techniques across various network structures and dynamic environments as tabulated in Table. 2. This includes traditional topology-based approaches (e.g., centrality and shortest paths), optimization and control-based frameworks, information diffusion models, machine learning-driven methods, and recent advances in dynamic and higher-order networks.
In this paper, we conduct a structured and comprehensive study of this field. We not only review the existing identification of key nodes in the network and the understanding of application scenarios from the perspective of network structure, methodology, and application scenario, but also discuss key challenges and unresolved issues. Note that the definitions of symbols used in this paper are provided in the Table 3.
Table 3: Table of Symbols.
Symbol Definition
$G$ Graph
$V=\{v_{1},v_{2},\dots,v_{n}\}$ Set of nodes
$E=\{e_{12},e_{13},\dots,e_{ij}|v_{i},v_{j}\in V\}$ Set of edges, $e_{ij}$ denotes an edge connecting nodes $v_{i}$ and $v_{j}$
$n=|V|$ Number of nodes
$m=|E|$ Number of edges
$\Gamma(v)$ Set of first-order neighbors of node $v$
$N(v)$ Set of first-order and second-order neighbors of node $v$
$d_{ij}$ Path between vertices $v_{i}$ and $v_{j}$
$\sigma_{ij}$ Shortest path length between two nodes
$A=\{a_{ij}\}_{n\times n}$ Adjacency matrix
$k(v_{i})$ Degree of vertex $v_{i}$
$\Delta_{ijk}$ Triangle formed by nodes $v_{i}$, $v_{j}$, and $v_{k}$
$X=[x_{1},x_{2},\dots,x_{n}]$ Eigenvector of the network
2 Neighbors-Based Ranking Methods
In complex network analysis, the contribution of neighbors can better measure the importance of nodes, and the essence of the node’s position and role in the network can be captured through domain information. This section discusses node ranking methods based on the neighbor information of nodes in the network topology as displayed in Table. 4. They can provide a detailed perspective on node importance by considering the local network structure, strike a balance between computational efficiency and accuracy, and are suitable for large-scale networks.
2.1 Degree Centrality
In the study of social network structures, Bavelas et al. RN2 introduced the concept of centrality, and Nieminen et al. RN3 addressed the complex and cumbersome calculation of point centrality metrics by proposing a direct measure of node importance through the adjacency matrix . Specifically, the degree of node $v_{i}$ is defined as follows:
$D^{\prime}(v_{i})=\sum_{j=1}^{n}a_{ij},\quad a_{ij}\in A=\{a_{ij}~\}_{n\times n}.$ (1)
The advantage of this metric is that it is intuitive and easy to understand, with simple computation. However, the node degree metric cannot have comparability across networks of different scales. Building on this, the concept of degree centrality for the entire network is introduced, defined as:
$k(v_{i})=\sum_{j=1}^{n}a_{ij}/n-1.$ (2)
The improved metric enables comparisons across networks of varying sizes. Intuitively, a node with more neighbors is likely to be more influential. However, this metric only considers the local state of node $v_{i}$ and cannot take into account the local environment of the node. To address the limitations of high computational cost and low local relevance in centrality measures, Chen et al. RN4 introduced semi-local centrality (SLC), which incorporates the degrees of both first and second-order neighbors with flexible parameters. While degree centrality performs effectively in scenarios with low diffusion probabilities, semi-local centrality is better suited for larger diffusion probabilities, capturing broader network influence. To mitigate the sensitivity of centrality measures to diffusion probabilitie. Ma et al.  ma2019quasi proposed Laplacian centrality, which measured node influence by the energy change of a network after node removal. It can be easily computed using node degree, significantly reducing computational complexity and offering good scalability, though it only considered nodes and their first-order information. Zhu et al.  zhu2024identifying introduced an Improved Laplacian Centrality (ILC) based on the self-consistency concept, incorporating global topological structure information.
2.2 K-shell Decomposition
Kitsak et al. RN41 proposed the $k_{s}$ index based on node local neighbors, which characterizes the topological structure of each node in the network. The method starts by removing all nodes with $k(v)=1$ from the network, and these nodes are defined as $k_{s}(v)=1$. This process repeats until only nodes with $k(v)\geq 2$ remain. Next, all nodes with $k(v)=2$ in the subgraph are removed, and these nodes are assigned an index of $k_{s}=2$. This process continues until only nodes with $k(v)\geq 3$ are left. The procedure stops when no nodes remain in the subgraph. In this way, each node is assigned to a corresponding $k$-shell layer, and the overall $k$-shell decomposition is obtained. This method has low time complexity, but it cannot distinguish the importance of nodes within the same layer.
To distinguish the importance of nodes within the same $k$-shell layer, the $k$-shell index overlooks the edges connected to the deleted nodes and the removal order of nodes during the node ranking process. Zeng et al. RN42 proposed the Mixed Degree Decomposition (MDD), which combines the degree of remaining and removed nodes to better distinguish nodes in the same $k$-shell layer. This method is effective in networks with special structures, such as trees, stars, and regular graphs. Li et al. li2018identification further classified neighbors by removal order and assigned weights to more accurately capture a node’s spreading capability. Liu et al. RN43 estimated node influence using the shortest distance to core nodes, while Zareie et al. zareie2018hierarchical used the distance to the network boundary as a proxy for core proximity. However, the $O(n^{2})$ computational complexity of shortest path calculations poses challenges for large-scale networks.
Extensive numerical simulations have shown that in some networks, nodes with high $k$-shell indices, even those classified as core nodes, are not always the most influential in information propagation RN43 ; zareie2018hierarchical . This indicates a discrepancy between the actual network core and the $k$-shell decomposition index. Ibnoulouafi et al. RN6 introduced M-Centrality, which integrates the $k$-shell index with the degree variation features of both nodes and their neighbors. This metric demonstrates robust performance in dynamic transmission and network connectivity, while maintaining a time complexity of $O(n)$. Liu et al. liu2015core proposed a link entropy metric that quantifies link diversity within shells by averaging the link diversity of nodes within the same shell. By normalizing for varying node counts, this method effectively distinguishes true core nodes from quasi-core groups. Additionally, Liu et al. liu2015improving argued that dense interconnections among core nodes could hinder effective information diffusion. They introduced edge diffusion importance by removing edges with low diffusion intensity between nodes, and applying $k$-shell decomposition to the pruned network, yielding a more accurate characterization of node importance.
2.3 H-index
H-index is originally designed to gauge the impact of academic journals and researchers. Lü et al. lu2016h extended the H-index to measure node importance in networks. The H-index of node $v_{i}$ is defined as:
$h_{i}=H(k_{j1},k_{j2},\dots,k_{jk_{i}}),$ (3)
where $H(\cdot)$ is the operator that finds the largest integer $h$ such that the node has at least $h$ neighbors with degree no less than $h$. If the H-index of node $v_{i}$ is $h$, it means that node $v_{i}$ has at least $h$ neighbors with degree no less than $h$. In large-scale social citation networks, nodes with the same H-index are often hard to distinguish. To address this, Liu et al. RN10 proposed the LH-index, incorporating both node and neighbor information. Gao et al. RN12 introduced the Weighted H-index for weighted networks by integrating edge weights. Furthermore, Zareie et al. zareie2019ehc proposed the Extended H-index centrality (EHC), which considers the cumulative contributions of all neighbors.
2.4 K-truss
Cohen et al. 2008Trusses introduced the concept of K-truss decomposition for a given unweighted undirected network. Let $e_{ij}$ represent the edge between nodes $v_{i}$ and $v_{j}$. In the K-truss decomposition, the support of edge $e_{ij}$ is the number of distinct triangles it forms, denoted as:
$sup(e_{ij})=\left|{\Delta_{ijk}:\Delta_{ijk}\in\Delta_{G}}\right|,$ (4)
where $\Delta_{ijk}$ represents a triangle with vertices $i$, $j$, and $k$, $k$ is a common neighbor of the two nodes, and $\Delta_{G}$ is the set of all triangles in network $G$. The higher the support of an edge, the higher its clustering characteristic. The decomposition proceeds based on edge support. Initially, set the parameter to $k=3$. Then, remove edges in the network with support less than k2k-2, and reduce the support of the two other edges forming a triangle with each edge by 1, updating the network topology. Continue removing edges with support less than k2k-2 until all edges in the network have support greater than or equal to k2k-2. Nodes removed in this stage are assigned to the k-truss layer. For the next step, increase $k+1$, and repeat the process of removing edges with support less than k2k-2 until all nodes in the network are completely assigned to the corresponding layers. Malliaros et al. malliaros2016locating introduced a triangle-graph-based k-core decomposition that refines the analysis of densely connected networks. Their method calculates the "trustness" of an edge, defined by its participation in a k-truss (the number of triangles it is part of). This approach offers a more detailed analysis compared to $k$-shell decomposition, making it particularly effective for identifying influential communities within networks.
2.5 Structural Hole
Yu et al. yu2017critical introduced the concept of "structural hole" to identify key nodes in a network by defining a node influence matrix $P$. If there is no direct connection or indirect redundancy relationship between two entities or communities, the obstruction between them is called a structural hole:
$p_{ij}=\frac{a_{ij}}{\sum_{j\in\Gamma(v_{i})}a_{ij}},$ (5)
where $p_{ij}$ represents the proportion of total effort that node $v_{i}$ contributes to maintaining node $v_{j}$ as a neighbor, and $\Gamma(v_{i})$ is the set of neighbors of node $v_{i}$. The network constraint coefficient of node $v_{i}$ is calculated as:
$C_{i}=\sum_{j\in\Gamma(v_{i})}\left(p_{ij}+\sum_{q}p_{iq}p_{qj}\right)^{2},\quad q\neq i,j,$ (6)
where $q$ is an intermediate node connecting $v_{i}$ and $v_{j}$, and $p_{iq}$ and $p_{qj}$ represent the total effort proportions of nodes $v_{i}$ and $v_{j}$ in maintaining their neighborhoods with a common neighbor $q$. A smaller constraint coefficient signifies a larger structural hole, making the node more influential in information propagation. Yu et al. yu2017identifying proposed an improved structural hole method (ISH) that integrates degree and nearest-neighbor information, effectively identifying critical nodes for information propagation by considering edge weights based on neighbor degrees.
In contrast, while the $k$-shell index prioritizes central nodes, it overlooks neighbor interactions and broader topological features. To address this, Xu et al. xu2018identifying introduced a local centrality measure (VKC) that combines neighborhood information with structural hole concepts, using outward-link diversity and network constraint coefficients. Building on these ideas, Liu et al. liu2021identifying54 developed an information control capability index (ICC), integrating $k$-shell values and enhanced structural hole theory. Zhao et al. zhao2023ranking21 further refined this approach with the SHKS node ranking index, combining $k$-shell and structural hole metrics, identifying vital bridging nodes with lower $k$-shell indices. While "structural holes" focus on nodes controlling parts of the network, they typically overlook the broader network topology. The combination of structural hole theory and $k$-shell methods enhances the overall network capacity ma2023effective .
2.6 Neighborhood Similarity
In a network, two similar nodes are more likely to be interested in transmitting the same information, and the likelihood of information propagating through the edges between the nodes is higher. Wang et al. wang2015link defined the similarity between nodes $v_{i}$ and $v_{j}$ based on the number of their common neighbors. Specifically, the more common neighbors two nodes have, the more similar they are. The similarity is mathematically defined as:
$HP(v_{i},v_{j})=\frac{\left|\Gamma(v_{i})\cap\Gamma(v(j))\right|}{\min(|\Gamma(v_{i})|,|\Gamma(v(j)|)},$ (7)
where $v_{i}$ and $v_{j}$ represent the starting node and the randomly selected neighboring node, respectively. Lu et al. lu2022critical11 proposed an Extended Similarity Coefficient Ranking Method (ESCRM), which connected the importance of nodes to the local structural similarity between nodes and their neighbors. Ai et al. ai2023identifying34 introduced Resource Allocation Similarity (RAS), simulating resource transfers based on the inverse sum of common neighbors to capture both structural and functional similarities. Rao et al. rao2022cbim decovered small communities from the multilayer network by dice neighborhood similarity (DNS), merging smaller communities and generating larger communities to maximize community influence and find k seed nodes in the multilayer network. Tong et al.  tong2024novel proposed an Extended Tanimoto Correlation (TC) for topological clustering of neighborhood sets to measure node importance.
Table 4: Comparison of Neighbors-Based Ranking Methods.
Title Related Works Advantages / Disadvantages
Degree Centrality RN2 RN3 RN4 ,  ma2019quasi zhu2024identifying ++ Simple and computationally efficient. - Fails to capture global network structure, only reflects local connectivity of neighbours.
$k$-shell Index RN41 RN42 ,  li2018identification ,  RN43 ,  zareie2018hierarchical ,  RN6 ,  liu2015core ,  liu2015improving ++ Suitable for large-scale networks, identifies core nodes. - Cannot differentiate nodes within the same layer, information spread may not correlate directly with $k$-shell values.
H-index lu2016h RN10 RN12 zareie2019ehc ++ Incorporates degree and neighbor influence. - Cannot distinguish nodes with the same H-index, dependent on neighborhood degree distribution.
K-truss 2008Trusses malliaros2016locating ++ Identifies community cores and modules. - High computational cost, ineffective in sparse networks.
Structural Hole yu2017critical yu2017identifying xu2018identifying liu2021identifying54 zhao2023ranking21 ma2023effective ++ Identifies bridging nodes, effective for information spread. - Computationally expensive, requires global information.
Neighborhood Similarity wang2015link lu2022critical11 ai2023identifying34 rao2022cbim tong2024novel ++ Beneficial for social networks and recommendation systems. - High computational complexity in large networks.
3 Eigenvector-Based Ranking Methods
Eigenvector-based methods evaluate node importance by considering both the number and influence of connections through eigenvector analysis and iterative computation as details in Table. 5. These methods extend beyond simple neighbor counting by incorporating the significance of connected nodes. They are extensively used in social network analysis to identify influential entities, in web search algorithms for page ranking, and in biological networks to detect critical components. The key strength of these methods lies in their ability to capture the recursive nature of importance: a node is deemed important if it is connected to other important nodes. This property enables more accurate assessment of node significance compared to approaches based solely on local neighborhood information. Nevertheless, such methods entail several limitations, including high computational cost, sensitivity to parameter settings, and the necessity for cautious interpretation across diverse network contexts.
3.1 Eigenvector Centrality
Eigenvector Centrality (EC) takes into account the importance of a node’s neighbors bonacich1971factoring  bonacich2007some . It uses the eigenvectors and eigenvalues of the network’s adjacency matrix to measure the importance of nodes. The main idea is to assume that the importance of node $v_{i}$ is denoted as $x_{i}$. Thus, the eigenvector centrality of node $v_{i}$ is defined as:
$EC(v_{i})=x_{i}=c\sum_{j=1}^{n}a_{ij}x_{j},$ (8)
where $c$ is the proportionality constant, and initializing $x=[x_{1},x_{2},\dots,x_{n}]$, after several iterations, it converges to $x=cAx$, rewritten as $Ax=c^{-1}x$, meaning $x$ is the eigenvector corresponding to the eigenvalue $c^{-1}$ of the adjacency matrix $A$. Building on this, Ilyas et al. ilyas2011identifying introduced Principal Component Centrality (PCC) to identify social hubs in large-scale networks, measuring node influence based on its position using eigenvectors of the adjacency matrix. To identify influential members in social networks, Estrada et al. PhysRevE71 defined subgraph centrality, which sums closed paths within a node’s subgraph, with shorter paths contributing more. Ahajjam et al. ahajjam2015leadersrank proposed the LeadersRank algorithm, ranking nodes based on eigenvector centrality and identifying communities through second-order neighbors. However, this algorithm primarily bases node importance on immediate neighbors, leading to slow convergence and neglect of external influences.
To address eigenvector centrality’s localization issue, which concentrates weight on a few nodes, Martin et al. martin2014localization introduced a method based on a non-backtracking matrix. This method avoids localization while yielding similar results to eigenvector centrality in dense networks. Zhong et al. zhong2018identifying combined Jaccard similarity with the dissimilarity between nodes and proposed the ECDS centrality method, which takes the weighted sum of the centralities of its neighboring nodes to form an iterative formula similar to the eigenvector centrality.
3.2 Katz Centrality
Katz centrality katz1953new extended eigenvector centrality by adding a constant term to describe the importance of the central node itself. It is defined as:
$x_{i}=c\sum_{j=1}^{N}A_{i,j}x_{j}+\beta,$ (9)
where $\beta$ is a constant, and in matrix form, it can be represented as $x=cAx+\beta$, which is equivalent to $x=(I-cA)^{-1}~\beta$. Here, $x$ represents the Katz centrality vector for all nodes in the network. It assigns different weights to all paths in the network, assuming that longer paths are less important than shorter ones. However, it has a high time complexity and is only applicable to smaller networks with few loops. To address the limitation of betweenness centrality, which only considers the shortest paths and ignores non-shortest paths, Zhang et al. zhang2015identifying proposed the BKC algorithm, which combines path-information and Katz centrality. This approach takes into account both local and global features, overcoming the limitation of evaluating node importance based solely on neighboring nodes.
3.3 Pagerank Algorithm
Traditional ranking methods for webpages rely on "keyword density, " but this can be undermined by "malicious keywords." Page et al. page1998pagerank proposed the core algorithm of Google’s search engine, Pagerank. The value of node $v_{i}$ at time $t$ is defined as:
$PR_{i}(t)=\sum_{j=1}{n}a_{ji}\frac{PR_{j}(t-1)}{k_{j}^{out}},$ (10)
where $k_{j}^{out}$ represents the out-degree of node $v_{j}$. However, in practice, there are cases where webpages are unreachable. To address this issue, a random jump probability $c$ is introduced:
$PR_{i}(t)=(1-c)\sum_{j}^{n}a_{ji}PR_{j}(t-1)/{k_{j}^{out}}+c/n,$ (11)
where $c$ is a constant derived from experiments. In the PageRank algorithm, each node has the same random jump probability, and the parameter $c$ can be flexibly adjusted based on the network characteristics. The algorithm measures the importance of a webpage based on the quality and quantity of its inbound links. If a webpage is linked by many pages or high-quality pages, it is considered to have greater importance.
To address the limitations of PageRank, such as its failure to incorporate node attributes or external information, Yang et al. yang2017mining described the frequency and duration of contact between nodes in a directed weighted network to capture the importance of nodes based on their in-degrees and out-degrees. They proposed the Two-Way-PageRank method to identify important nodes in directed weighted networks. Hsu et al. hsu2017unsupervised introduced the AttriRank model, which integrates node attribute similarity into the ranking process. This model utilizes random walks, capturing associations between nodes based on both network structure and attribute similarity. The random walk is framed as a Markov chain, with a transition matrix that balances information propagation across the network and node attributes. Building on this, Sheng et al. sheng2020identifying12 proposed the Trust-PageRank (TPR) algorithm, which combines attribute similarity and trust values, defined through structural degree ratios. TPR effectively merges node attributes and network structure to refine node rankings. Su et al. su2021identification67 introduced a cascade failure and directionally weighted PageRank method tailored for power grid networks. This approach incorporates grid connectivity, electrical topology, and power flow direction, thus enhancing the grid’s resilience by mitigating the impact of cascading failures.
Table 5: Comparison of Eigenvector-Based Ranking Methods.
Title Related Works Advantages / Disadvantages
Eigenvector Centrality bonacich1971factoring bonacich2007some ilyas2011identifying ,  PhysRevE71 ,  ahajjam2015leadersrank martin2014localization zhong2018identifying ++ Considers neighbor importance, widely used in networks. - Suffers from localization issues, slow convergence.
Katz Centrality katz1953new zhang2015identifying ++ Balances local and global influence. - High computational cost, sensitive to parameters.
PageRank Algorithm page1998pagerank yang2017mining hsu2017unsupervised ,  sheng2020identifying12 su2021identification67 ++ Scales well for large networks, robust. - Sensitive to damping factor, ignores node attributes.
Voting Mechanism RN33 ,  li2014identifying ,  RN36 RN38 ,  RN37 ,  li2022improved23 ,  liu2016evaluating ++ Simple and adaptable for various networks. - Affected by network structure, struggles in dense graphs.
3.4 Voting Mechanism
To address the parameter adjustment issue in the PageRank algorithm, Lü et al. RN33 introduced a background node $g$ in the network, which establishes bidirectional connections with all other nodes, making the network strongly connected. They proposed the LeaderRank (LR) algorithm. Initially, all pages are assumed to have a LR value of 1, while the background node has a LR value of 0. After convergence, the LR value for node $i$ is defined as:
$LR_{i}=LR_{i}(t_{c})+\frac{LR_{g}(t_{c})}{n}.$ (12)
After traversing all the nodes in the network, the LR value is obtained. Compared to PageRank, LeaderRank has stronger robustness in resisting spammer attacks and random interference. Building on this, Li et al. li2014identifying introduced the weight-LeaderRank algorithm for weighted networks, which modifies the standard random walk by incorporating a biased random walk. This adjustment ensures that neighboring nodes receive higher scores, thereby improving the identification of influential propagators. To address the issue of overlapping influence in methods such as PageRank, ClusterRank, and $k$-shell, Zhang et al. RN36 proposed the iterative VoteRank method. In this approach, nodes receive votes from their neighbors, and their importance is determined by the number of votes they accumulate. This method reduces to degree-based ranking when all nodes have the same voting ability.
Building on the information of node neighborhood within the network topology, Kumar et al. RN38 proposed the NCVoteRank method, which adjusts the voting weights according to the core degree of neighboring nodes, adding flexibility to the voting framework. Sun et al. RN37 enhanced this concept by introducing the weight-VoteRank method, which incorporates both the number of neighbors and edge weights, allowing for a more nuanced measure of node importance in weighted networks. Li et al. li2022improved23 further refined the VoteRank algorithm with the introduction of Linearity and Importance (DIL) liu2016evaluating . This method initializes node scores and voting abilities based on local importance, and updates the voting capabilities of neighbors in each round. Through iterative voting, the Top-$k$ nodes with the highest scores are selected, improving the efficiency and accuracy of the ranking process. These methods collectively advance the VoteRank algorithm by integrating node neighborhood information, edge weights, and local importance, thereby enhancing the precision and efficiency of node ranking in complex networks.
4 Path-Based Ranking Methods
Path-based ranking methods assess node importance through the analysis of paths connecting nodes, taking into account their positions and roles within network connectivity and information dissemination. These methods characterize the global structure and connection patterns of networks by evaluating shortest paths, average distances, and random walk processes as details in Table. 6. Applications include identifying intermediaries and pivotal communicators in social networks, locating critical hubs in transportation systems, and understanding material or information transfer mechanisms in biological networks. The primary advantage of these approaches lies in their capability to capture the global influence of nodes and highlight those that function as bridges or central connection points within the network. However, these methods are computationally intensive, particularly for large-scale networks, necessitating simplifications or approximations to ensure practical feasibility.
4.1 Betweenness Centrality
Freeman et al. RN56 reviewed traditional centrality measures to determine the importance of nodes in social networks. Based on the connectivity between nodes, they introduced Betweenness Centrality (BC), which measures the importance of a node by counting the number of shortest paths that pass through the node. It is defined as:
$BC(v)=\sum_{i,j\neq v}~\frac{\left|d_{ij}(v)\right|}{\left|d_{ij}\right|},$ (13)
where $\left|d_{ij}\right|$ is the number of shortest paths between nodes $v_{i}$ and $v_{j}$, and $\left|d_{ij}(v)\right|$ is the number of shortest paths between nodes $v_{i}$ and $v_{j}$ that pass through node $v$. BC takes into account the global features of the network, but its path computation has a high time complexity. Hage et al. hage1995eccentricity proposed using the maximum distance between a node and all other nodes in the network as a measure of the node’s eccentricity centrality, ECC. The eccentricity centrality of node $v_{i}$ is defined as:
$\text{ECC}(v_{i})=\max_{j}(d_{ij}),\quad j=1,2,\dots,n,$ (14)
where the network diameter is the maximum eccentricity value among all nodes in the network $G$, and the network radius is defined as the minimum eccentricity value among all nodes. Clearly, the node whose eccentricity centrality value equals the network radius is the central node. Boccaletti et al. RN59 proposed the ASP index, which measures a node’s importance based on its effective information transfer to all other nodes. The ASP index is defined as the sum of the shortest path distances from a node to all other nodes, standardized by the total number of node pairs in the network. However, the ASP index becomes ineffective when a node is removed, as it can disconnect the network.To address this, Lü et al. RN61 introduced the Relative Change of Average Shortest Path (RASP) centrality, which focuses on the change in the average shortest path in the network after a node is removed. RASP measures the node’s importance in maintaining network connectivity by calculating the relative change in the average shortest path.
To address the high time complexity of betweenness centrality, Song et al. song2015novel introduced load centrality (LC), which measures the change in network paths after node removal to assess the node’s impact on connectivity. Ventresca et al. ventresca2015efficiently developed the CNICC algorithm, which determines node connectivity and capacity based on path length, the number of paths, and network scale. Many real-world networks display hierarchical and modular structures. Zhang et al. zhang2011node proposed a multi-scale node importance metric using a kernel function, where bandwidth determines the range of interactions. Small bandwidth captures short-range interactions, while large bandwidth highlights long-range interactions, offering insights into node influence across different scales during dynamic processes. In power systems, Yang et al. yang2020critical introduced the Electric Observability Capability Index (EOCI) and the Electrical Dynamic Characteristics Index (EkI) to identify critical nodes influencing power system observability and controllability. Kianian et al. kianian2021efficient used degrees and independent influence paths to approximate influence propagation, reducing computational cost by pruning insignificant nodes. Xiao et al. xiao2024new applied LASPN theory to identify critical nodes, combining local average shortest paths and extended neighborhoods for efficient processing in large-scale networks.
4.2 Closeness Centrality
To address the interference of special values in paths, Freeman et al. freeman2002centrality introduced Closeness Centrality by calculating the average distance of a node to all other nodes in the network. The smaller the average shortest distance $d_{i}$ from node $v_{i}$ to other nodes in the network, the closer node $v_{i}$ is to the other nodes in the network. The reciprocal of $d_{i}$ is defined as the closeness centrality of node $v_{i}$:
$CC(i)=\frac{1}{d_{i}}=\frac{n-1}{\sum_{j\neq i}d_{ij}}.$ (15)
In a connected network, the smaller the average distance between a node and all other nodes in the network, the larger the closeness centrality of that node. This can be understood as determining the importance of a node by the average transmission time of information in the network. To improve closeness centrality for network efficiency, Latora et al. latora2001efficient extended betweenness centrality to disconnected networks and proposed the EFF index. However, this approach requires calculating the distances between all node pairs, which is computationally expensive for large networks. To reduce this complexity, Salavati et al. salavati2019ranking applied the Louvain community detection algorithm to identify network communities. By maximizing modularity, they extracted community structures and selected key nodes using betweenness centrality. They also introduced a method to identify gateway nodes linking communities and reduced computation by focusing on a subset of special nodes. Okamoto et al. Okamoto7328 proposed the RAND algorithm to estimate average node distances efficiently. It selects top candidate nodes based on these estimates and ranks them precisely, balancing computational efficiency with accuracy. Sheng et al. sheng2020identifying introduced the GLS method, which calculates the influence of nodes by considering common neighbors and structural characteristics. This method combines local and global structural influence based on their information exchange capabilities.
Table 6: Comparison of Path-Based Ranking Methods.
Title Related Works Advantages / Disadvantages
Betweenness Centrality RN56 ,  hage1995eccentricity ,  RN59 ,  RN61 ,  song2015novel ,  ventresca2015efficiently ,  zhang2011node ,  yang2020critical ,  kianian2021efficient ,  xiao2024new ++ Captures shortest paths, global influence, robustness. - High computational cost, sensitive to network changes.
Closeness Centrality freeman2002centrality latora2001efficient salavati2019ranking Okamoto7328 sheng2020identifying ++ Measures reachability efficiency, effectiveness. - Ineffective in disconnected networks, computationally expensive in large graphs.
Random Walk iannelli2018influencers RN7 kermarrec2011second ++ Models information spread, considers network topology, improves influence ranking. - Computationally intensive, rely on walk length.
4.3 Random Walk
Currently, methods for identifying key nodes based on heuristic centrality measures are used only for specific network topologies or unique dynamic models for analysis and demonstration. Iannelli et al. iannelli2018influencers proposed a method based on the random walk effective distance $D^{RW}_{ij}(\lambda)$ between pairs of nodes, using the adjacency matrix $A_{ij}$ and the diffusion dynamic parameter $\lambda$ to quantify the influence of nodes during the network diffusion process. The ViralRank score for node $v_{i}$ is defined as the average random walk effective distance between all source nodes and target nodes in the network:
$v_{i}(\lambda)=\frac{1}{N}\sum_{j}\left(D_{ij}^{RW}(\lambda)+D_{ji}^{RW}(\lambda)\right),$ (16)
$D_{ij}^{RW}(\lambda)=-\text{In}~\sum_{k\neq j}\left(I^{j}-e^{-\lambda}P^{(j)}\right)^{-1}_{ik}e^{-\lambda}p_{k}^{(j)},$
where $i\neq j$, $D_{ii}^{RW}=0$, $P^{(j)},I^{(j)}\in R^{(N-1)\times(N-1)}$ are partition matrices of the Markov matrix, $(P)_{ij}=A_{ij}/\sum_{k}A_{ik}$, and the identity matrix $(I)_{ij}=\theta_{ij}$ are obtained by removing the $i$-th row and $j$-th column, respectively. $p^{(j)}$ is the $j$-th row of $P$ after removing the $v_{j}$-th element. A node’s smaller average effective distance means a higher ViralRank score, indicating central nodes cause larger outbreaks than peripheral ones as "seed" nodes. Dong et al. RN7 proposed the Semi-Local Centrality (SLC) method, integrating topological features with dynamic processes. SLC starts a random walk from node $v_{i}$, adding influential neighboring nodes to the path. The semi-local centrality of a node is defined by the length of the random walk path formed by its influential neighbors, improving the accuracy. To address the limitations of current centrality measures and their incompatibility with distributed environments, Kermarrec et al. kermarrec2011second proposed second-order centrality, which calculates the standard deviation of return times recorded by each node during a random walk. This approach supports distributed computation and provides a local value for each node, reflecting its relative importance. It allows unbiased identification of key nodes through random walk, enabling a global understanding of network topology.
5 Control-Optimization Based Ranking Methods
Control optimization-based ranking methods identify key nodes by employing network control theory and optimization techniques. These methods aim to characterize the influence of nodes on network dynamics, controllability, and connectivity, offering valuable insights for applications in security, robustness enhancement, and regulation of dynamic processes as details in Table. 7. In security, they help detect vulnerable links and improve resilience against attacks. In infrastructure networks, they aid in designing effective control strategies. The primary advantage of these approaches lies in their ability to offer a dynamic and interactive perspective on node importance, revealing the impact of individual nodes on overall network functionality and stability. However, these methods face challenges such as high computational complexity as well as control and optimization theory.
5.1 Network Control
Existing research has highlighted the importance of specific nodes in regulating network dynamics and functionality. Li et al. li2006controlling employed the Watts-Strogatz propagation model to explore spreading mechanisms in small-world networks. Their analysis of delayed control propagation with linear and nonlinear feedback controllers, demonstrating how network parameters influence stability and oscillatory behavior. Ghosh et al. ghosh2022synchronized emphasized the critical role of dynamic synchronization in maintaining global network coherence. Their findings showed that proper weighting procedures significantly enhance synchronization in static complex networks, with key nodes exerting disproportionate influence on system stability. D’Souza et al. d2023controlling proposed a framework for network control that integrates node dynamics and macroscopic network properties, refining control strategies through higher-order interactions. Liu et al. liu2011controllability developed analytical tools to assess the controllability of arbitrary directed networks, establishing that the number of driver nodes required for full control is primarily dictated by the network’s degree distribution. Their work uncovered a fundamental relationship between network structure and control feasibility: sparse, heterogeneous networks are the most difficult to control, whereas dense, homogeneous networks require fewer driver nodes. The controllability of a system governed by linear time-invariant dynamics:
$\frac{dx(t)}{dt}=Ax(t)+Bu(t),$ (17)
where $x(t)\in R^{N}$ represents the system state of $N$ nodes at time $t$, $A\in R^{N\times N}$ is the adjacency matrix encoding interaction dynamics, and $B\in R^{N\times M}$ identifies the driver nodes influenced by control inputs $u(t)\in R^{M}$. The system is controllable if and only if the controllability matrix, $C=(B,AB,\dots,A^{n-1}B)\in\mathbb{R}^{N\times NM}$. The system is controllable if and only if $Rank(C)=N$, meaning it can transition from any initial state to any final state within a finite time. Furthermore, the minimum number of driver nodes $N_{D}$ required for full control can be estimated using the structural controllability framework as $N_{D}=\max\{1,N-\operatorname{rank}(A)\}.$ To address the problem of identifying an optimal driver node set for minimizing control cost, recent studies have proposed formulations such as
$\min_{B}|u(t)|^{2}\quad\text{subject to}\quad\frac{dx(t)}{dt}=Ax(t)+Bu(t),$ (18)
which seeks the control input $u(t)$ with the smallest energy expenditure required to guide the system toward the desired state. To address the minimal-cost control problem, Ding et al. ding2017key proposed the Projection Gradient Method Extension (R-PGME) algorithm, which leverages Monte Carlo simulations and identifies a minimal set of key nodes for effective control. Lu et al. lu2019pinning applied the Warshall algorithm to Boolean control networks (BCNs), determining the smallest set of nodes required for stability in large-scale systems. Expanding on these approaches, Zhu et al. Zhu10106394 integrated structural information and developed a low-complexity structural controller for large-scale Boolean networks, efficiently identifying the minimal feedback node set for robust control.
The eigenvalues of the Laplacian matrix play a critical role in pinning control and network controllability, directly influencing the selection of optimal driver nodes for synchronizing large-scale networks. Yu et al. yu2012distributed proposed a distributed adaptive strategy that adjusts coupling weights based on local node dynamics to achieve synchronization, later extending their work to establish synchronization criteria for various network topologies, including strongly and weakly connected graphs yu2013synchronization . Ding et al. ding2016optimizing demonstrated that denser and more homogeneous networks exhibit superior controllability, with controlled nodes typically having lower in-degree than the network average. Amani et al. amani2017finding introduced the Eigenvalue Sensitivity Index (ESI), which ranks nodes using extremal Laplacian eigenvectors to estimate their importance in pinning control. They further developed a computationally efficient method based on sensitivity analysis of the Laplacian matrix to identify near-optimal driver nodes amani2018controllability . Liu et al. liu2018optimizing evaluated pinning effectiveness via the smallest eigenvalue of the grounded Laplacian, linking controllability to network structure. Several optimization strategies have been investigated. Wang et al. wang2012optimizing achieved full network control through structural perturbations using a single driver node. Bof et al. bof2016role showed that energy-efficient control depends on the left and right Perron eigenvectors of the network matrix. Zhou et al. zhou2018node proposed the ControlRank metric, validating degree-based rankings in scale-free networks. Liu et al. hui2021node refined driver node selection by prioritizing nodes with the largest eigenvector components corresponding to the dominant Laplacian eigenvalue. Bomela et al. bomela2023finding applied the Moore–Penrose pseudoinverse of the Laplacian to identify Most Influential Nodes (MIN) in oscillatory networks. Recent studies have emphasized minimal intervention strategies. Jiang et al. jiang2024pinning showed that large-scale synchronization can often be achieved with minimal control effort or higher-order interactions. Sun et al. sun2025optimal analyzed optimal control in reaction-diffusion networks, revealing the influence of topology and diffusion parameters on epidemic dynamics. Collectively, these works highlight the strong interdependence between spectral properties, network structure, and dynamic processes, underscoring the value of targeted interventions for achieving efficient and robust network control.
The contribution of network edges, including loops, to overall network dynamics is fundamentally governed by the eigenvector corresponding to the smallest nonzero eigenvalue of the Laplacian matrix, known as the Fiedler vector. The cumulative effect of all edges is encapsulated in the network’s algebraic connectivity, which dictates synchronization and diffusion properties. Zhang et al. zhang2017effect analyzed the impact of adding feedback edges in directed acyclic graphs (DAGs) on consensus behavior in multi-agent systems. They identified disruptive feedback edges that degrade convergence speed and formulated a necessary and sufficient condition for their identification. Mo et al. mo2019effects extended this by introducing a topology-based optimization framework for directed graphs, enabling the systematic selection of edges that enhance convergence. They demonstrated that in strongly connected weighted digraphs, the second smallest Laplacian eigenvalue remains real and established conditions for edge selection that improve consensus dynamics. Jiang et al. jiang2023searching proposed a method to assess the importance of loops in complex networks, ranking key loops based on their influence on the Fiedler value. Similarly, Cao et al. cao2024synchronization developed a theoretical framework for synchronization acceleration by optimizing directed edge placement and weights. They derived a necessary and sufficient condition for selecting edges that enhance convergence, allowing efficient identification through Laplacian eigenspace computations. Zhang et al. zhang2022necessary reaffirmed the existence of disruptive feedback edges in DAGs and explicitly constructed conditions for their identification. Gao et al. gao2025effects further examined how weighted edge additions influence algebraic connectivity in directed graphs, proving that changes in Laplacian eigenvalues are localized within subgraphs containing the added edges. Their findings underscore the dependence of algebraic connectivity on edge weight, range, and distribution along network paths. These studies further reveal how interventions of network edges and cycles control dynamic behavior, highlighting the role of spectral properties in optimizing synchronization and controllability.
5.2 Maximization of Connected Component
In a network, there exists a set of important nodes that influence the network’s robustness and connectivity. By removing these nodes or attacking the network, we can identify a critical set of nodes. Deleting this set of nodes can minimize or maximize the connectivity components (such as the largest connected component or node pair connectivity) in the remaining network, which is known as the Critical Nodes Deletion Problem. Aringhieri et al. aringhieri2016general designed a general framework to address several classic critical node problems (CNPs), specifically: (1) Pairwise connectivity: Minimize the number of connected node pairs in the remaining network by deleting up to $K$ nodes, ensuring the pairwise connectivity is reduced to a threshold $P$. (2) Largest connected component: Minimize the size of the largest connected component in the remaining network by deleting up to $K$ nodes, ensuring its size cannot exceed a threshold $L$. (3) Number of connected components: Maximize the number of connected components in the remaining network by deleting the fewest nodes possible, ensuring it meets or exceeds a threshold $N$. It discusses greedy algorithm design rules for these problems and proposes general algorithms for solving various CNPs, providing solutions for subsequent derivations and improvements of CNP models. Alozie et al. alozie2021efficient focused on optimizing distance-based connectivity metrics by removing a subset of nodes. They aimed to minimize two specific metrics: (1) the number of node pairs connected by the longest path, and (2) the Harary index or efficiency of the graph. To achieve this, they proposed a general mixed-integer programming formulation and a separation algorithm based on breadth-first search tree generation, designed to solve the constraint inequalities.
Thulasiraman et al. Thulasiraman evaluated node value using cut vertices and edges for network diagnostics and security, employing link testing and node connection statistics. In protein interaction networks, Boginski et al. Boginski0007 aimed to identify key nodes by solving the Cardinally Constrained-CNDP (CC-CNDP), defined as minimizing the set of deleted nodes $R$ such that the largest connected component in the remaining subgraph $G(V\setminus A)$ of the network cannot exceed $L$, as follows:
$\text{minimize}\quad\sum_{i\in V}v_{i}$ (19)
s.t. $e_{ij}+v_{i}+v_{j}\geq 1,\quad\forall(i,j)\in E,\quad(1)$
$e_{ij}+e_{jk}-e_{ki}\leq 1,\quad\forall(i,j,k)\in V,\quad(2)$
$e_{ij}-e_{jk}+e_{ki}\leq 1,\quad\forall(i,j,k)\in V,\quad(3)$
$-e_{ij}+e_{jk}+e_{ki}\leq 1,\quad\forall(i,j,k)\in V,\quad(4)$
$\sum_{i,j\in V}e_{ij}\leq L,\quad(5)$
$v_{i}\in\{0,1\},\quad\forall i\in V,e_{ij}\in\{0,1\},\quad\forall i,j\in V,\quad(6)$
where $e_{ij}$ represents the edge between nodes $v_{i}$ and $v_{j}$, (1) ensures that there is an edge between nodes $v_{i}$ and $v_{j}$, meaning that both nodes are in the network, (2-4) constrain the connectivity of the node set $(i,j,k)$ to be minimal, and (5) constrains the size of the largest connected component in the network. This CNDP problem is transformed into an integer programming optimization problem to identify critical network disruptions and target proteins for drug neutralization. In self-organizing networks, Karygiannis et al. Karygiannis1644271 defined critical nodes as breakpoints under failures or attacks. They used graph theory’s vertex and edge cut principles to reveal network vulnerability through subgraph generation and involved solving optimization problems where their removal satisfies specific constraints.
Table 7: Comparison of Control Optimization-Based Ranking Methods.
Title Related Works Advantages / Disadvantages
Network Control li2006controlling ,  ghosh2022synchronized ,  d2023controlling ,  liu2011controllability ,  ding2017key ,  lu2019pinning ,  Zhu10106394 ,  yu2012distributed ,  yu2013synchronization ,  ding2016optimizing ,  amani2017finding ,  amani2018controllability ,  liu2018optimizing ,  wang2012optimizing ,  bof2016role ,  zhou2018node ,  hui2021node ,  bomela2023finding ,  jiang2024pinning ,  sun2025optimal ,  zhang2017effect ,  mo2019effects ,  jiang2023searching ,  cao2024synchronization ,  zhang2022necessary ,  gao2025effects ++ Beneficial for dynamic networks and control applications, quantifies node controllability. - Requires accurate network topology information, relies on control theory.
Maximization of Connected Component Optimization aringhieri2016general ,  alozie2021efficient ,  Thulasiraman ,  Boginski0007 ,  Karygiannis1644271 ,  arulselvan2011cardinality ,  li2011finding ,  ventresca2015experimental ,  veremyev2014exact ,  ren2020identifying ,  wang2010community ,  lam2013identifying ,  ventresca2014fast ,  tang2020discrete ++ Increases network fragmentation, applicable to security and IDS applications. - Excessively breaks the network, affecting normal operations.
Minimization of Pair Connectivity Optimization di2012branch ,  ARULSELVAN20092193 ,  shen2012adaptive ,  ventresca2014derandomized ,  dinh2015assessing ,  sarker2019critical ,  purevsuren2016heuristic ,  addis2016hybrid ,  chen2020critical ,  ventresca2012global ,  shen2012discovery ,  ventresca2014region ,  yin2023mixed23 ,  zhang2024evolutionary ,  fortz2024min ,  jiang2025identifying ,  kouam2025novel ,  zhou2024finding ++ Effective in suppressing spreading and network attacks. - High computational demand in large-scale network.
Minimum Spanning Tree chen2004evaluation ,  di2011complexity ,  hermelin2016parameterized ,  addis2013identifying ,  aringhieri2016local ,  wang2017evaluation ++ Simple and computationally efficient. - cannot fully capture network complexity and dynamics.
Furthermore, Arulselvan et al. arulselvan2011cardinality proposed a risk management-based method to detect a set of nodes that, when deleted, cause the connectivity index of the resulting subgraph to fall below a threshold. This problem is defined as the cardinality-constrained critical node problem (CC-CNP) and modeled as an integer linear programming problem, which is solved using a heuristic genetic algorithm. In social networks, there exists a set of $k$ mediator nodes aimed at maximizing the probability of information activation from a source node to a target node. Li et al. li2011finding defined this problem as the k-Mediators problem. Based on the independent cascade propagation model, they optimized information propagation by considering path probability effects and structural connectivity, using the product of edge weights to more accurately simulate the information spread process. They proposed a three-step greedy method to iteratively construct the best propagation tree (BPT), connecting all source and target nodes and optimizing the propagation paths to maximize activation probabilities. Ventresca et al. ventresca2015experimental developed a multi-objective critical node detection algorithm. It removes nodes to maximize the connected components in the remaining network while minimizing the size variance among these components. Veremyev et al. veremyev2014exact introduced two critical node deletion problems: (1) deleting $K$ nodes to minimize the remaining connected node pairs; (2) deleting the fewest nodes to ensure the largest component has no more than $L$ nodes. They reduced the solving time through constraints, making it possible to tackle large-scale network optimization issues. Ren et al. ren2020identifying proposed a reverse greedy strategy to identify nodes crucial for maintaining connectivity. Starting from an empty network, it adds nodes with the least impact on the largest component, ranking by addition order, with the last added node deemed most critical.
To address the high computational costs of greedy algorithms on large networks. Wang et al. wang2010community developed a community-based greedy algorithm using node edge weights for propagation speed, incorporating information diffusion for community detection and dynamic programming to identify influential nodes. Lam et al. lam2013identifying used simulations in the NetLogo multi-agent environment to identify critical infrastructure components, assessing node failure impacts via attack scenarios and the giant component metric. Ventresca et al. ventresca2014fast proposed a depth-first search method with $O(|V|+|E|)$ complexity, integrated into greedy algorithms for quick key node identification in large networks, enabling efficient CNDP objective function solving. Tang et al. tang2020discrete introduced a Discrete Shuffling Frog Leaping Algorithm based on virtual frog population topology, featuring novel encoding, evolution rules, and a local degree-based replacement strategy for efficiency.
5.3 Minimization of Pair Connectivity
To minimize pairwise connectivity by removing a fixed number of nodes, Di Summa et al. di2012branch proposed an integer linear programming (ILP) model with non-polynomial constraints, as follows:
$\text{maximize}\quad\sum_{i,j\in V,i\neq j}e_{ij}$ (20)
s.t. $\sum_{i\in V}v_{i}\leq K,\quad(1)$
$v_{i}+v_{j}\geq e_{ij},i,j\in E,\quad(2)$
$e_{ij}+e_{jk}-e_{ik}\geq 0,i,j,k\in K,i<j<k,\quad(3)$
$e_{ij}-e_{jk}+e_{ik}\geq 0,i,j,k\in V,i<j<k,\quad(4)$
$-e_{ij}+e_{jk}+e_{ik}\geq 0,i<j<k,\quad(5)$
$v_{i}\in\{0,1\},e_{ij}~\in\{0,1\},i,j\in V,i<j,\quad(6)$
where constraint (1) limits the number of deletions to $K$, while (2) ensures that removing an edge implies removing both its endpoints. Constraints (3)–(5) enforce connectivity dependencies, ensuring that if $\{i,j\}$ is disconnected in the remaining graph, then the node pairs $(i,k)$ and $(k,j)$ cannot be connected.
In small-scale networks, heuristic algorithms are commonly used to identify critical nodes that impact connectivity. Arulselvan et al. ARULSELVAN20092193 employed a combinatorial heuristic integrating integer programming to efficiently minimize pairwise connectivity by selectively removing nodes. Similarly, Shen et al. shen2012adaptive introduced an adaptive algorithm that identifies critical nodes without requiring full recalculations after topological changes, enhancing efficiency and response speed. Ventresca et al. ventresca2014derandomized developed an approximation algorithm for undirected, unweighted graphs, combining integer linear programming with randomized rounding. By handling vertex and triangle inequality constraints separately, their method effectively reduces subgraph connectivity while capping the number of deleted nodes at $k$, with solution quality bounded by a constant factor. Dinh et al. dinh2015assessing assessed network vulnerability under dynamic conditions using expected pairwise connectivity (EPC) as a metric, transforming the problem into a stochastic optimization framework based on probabilistic graphs. Sarker et al. sarker2019critical proposed an enumeration algorithm that accounts for connection costs and node weights, achieving high time efficiency and controlled complexity. Their approach, applied to river drainage networks, confirms power-law fragmentation behavior following the removal of critical nodes.
To address the complexity and accuracy balance in identifying critical nodes in large-scale networks, Pullan et al. purevsuren2016heuristic proposed a Greedy Randomized Adaptive Search Procedure (GRASP) with Path Relinking. The algorithm introduces an evolutionary path relinking mechanism during the path search process. It was shown that the GRASP with Path Relinking is more efficient than algorithms like neighborhood search and simulated annealing in solving the Critical Node Deletion Problem. Addis et al. addis2016hybrid proposed a "hybrid" heuristic algorithm based on the basic greedy algorithm. By combining two greedy rules, "node addition and deletion, " the algorithm addresses the CNDP problem of removing K nodes to minimize the residual network connectivity. The algorithm alternates between node addition and deletion operations in the feasible solution space, effectively avoiding local optimal solutions. Chen et al. chen2020critical focused on the critical node fragmentation problem in weighted networks. They proposed a non-convex mixed-integer quadratic programming (MIQP) model for undirected weighted networks. A greedy algorithm is employed to approximate the optimal solution, and the algorithm iteratively selects K critical nodes (Critical Node Set, CNS) that cause the most significant decrease in network pairwise connectivity, identifying the critical node combination that maximizes network performance.
To mitigate the high computational complexity of the critical node deletion problem, Ventresca et al. ventresca2012global employed a population-based incremental learning algorithm combined with combinatorial simulated annealing to detect critical nodes, leveraging depth-first search to minimize pairwise connectivity. For dense networks where pairwise connectivity computations remain $O(n^{2})$, Shen et al. shen2012discovery developed a hybrid iterative linear programming rounding (HILPR) algorithm, which transforms the problem into ILP, improving solving efficiency through local search and constraint pruning. Ventresca et al. ventresca2014region proposed a region-growing dual-criterion method using LP relaxation to iteratively expand a fixed-radius ball, providing a logarithmic approximation for optimal node removal. Addressing weighted social networks, Yin et al. yin2023mixed23 introduced the Mixed Strength Decomposition (MSD) method, dynamically setting thresholds for node removal based on connection weights, improving the precision of critical node identification. These advancements enhance computational efficiency in solving CNDP, enabling more effective identification of critical nodes in large-scale networks.
Recent studies have focused on optimizing critical node detection and network robustness under adversarial conditions. Zhang et al. zhang2024evolutionary formulated the Critical Node Detection problem in interdependent networks as a Bi-objective Optimization Problem (BICND) and designed a method integrating single-layer and multi-layer network transfer. This approach reduced algorithmic complexity while enhancing accuracy in identifying critical nodes. Fortz et al. fortz2024min proposed a compact integer programming formulation for the node attack optimization problem using pseudo-components within a two-level model, addressing performance limitations of prior iterative methods. Jiang et al. jiang2025identifying introduced the EBC method, leveraging loop structures to assess network robustness against targeted attacks and precisely identify edges critical to resilience. Kouam et al. kouam2025novel simulated adversarial behavior to evaluate node influence, modeling attack progression toward the target. Zhou et al. zhou2024finding examined the NP-hard nature of the protection node problem and optimized key node selection via an influence redundancy mechanism. By minimizing the adjacency matrix’s spectral radius after key node removal, they demonstrated submodular properties in the objective function and achieved an approximation ratio of $O(N\log N)$. These advances enhance network resilience by refining critical node and edge, offering efficient solutions for targeted attacks. The multi-area power system can be described using power grid models and associated control models. In contrast, the networked TST-S fuzzy system is modeled using TST-S fuzzy models that consist of multiple fuzzy rules. The research on "Critical-metrics-based attack strategy and resilient $H_{\infty}$ state estimator design for multi-area power systems" presents a critical-metrics-based attack strategy. From the attacker’s perspective, this strategy targets the critical information packets in the key areas of the power system, thereby amplifying the impact of the attack on system performance. Correspondingly, from the defender’s perspective, a resilient state estimator is designed to alleviate the adverse effects of potential critical-metrics-based attacks.
5.4 Minimum Spanning Tree
Chen et al. chen2004evaluation evaluated the relative importance of two sets of nodes by comparing the number of spanning trees in the remaining network after node deletion. The fewer the number of spanning trees in the corresponding network after removing a set of nodes and their associated edges, the more important that set of nodes is considered to be. This method is simple to compute and provides a more accurate reflection of node importance. When the network size is large, the algorithm complexity of solving the problem increases. To improve the solving efficiency, the objective function is decomposed into multiple subproblems. Di Summa et al. di2011complexity studied the minimization of the number of connected node pairs in the remaining graph after node deletion, designing the network as a tree structure to extend the solution to the CNP subproblem. Hermelin et al. hermelin2016parameterized discussed the Critical Node Cut (CNC) problem: given a graph $G$ with $n$ vertices, determine $k$ nodes, and after their removal, the remaining graph has at most $x$ connected node pairs. Using a parameterized method, they proved that the problem can be solved in polynomial time by preprocessing it into a subproblem with lower complexity.
Regarding the linear framework of the network traffic and connectivity constraint problem, Addis et al. addis2013identifying focused on the combinatorial problem of node deletion. They provided a dynamic programming recursive method for solving the tree decomposition problem based on graphs, which can solve the CNP problem in polynomial time on general and special graphs (split graphs, bipartite graphs, and the complements of bipartite graphs). Aringhieri et al. aringhieri2016local proposed two metaheuristics based on iterative local search and variable neighborhood search frameworks to achieve maximum fragmentation of graphs. They designed two computationally efficient neighborhoods, evaluated different exploration strategies, and found that the improved strategy produced a large number of new best solutions. Using betweenness centrality, they were able to reduce the runtime without compromising the solution quality. Wang et al. wang2017evaluation combined the advantages of the minimum spanning tree (MST) structure and the minimum connected dominating set (MCDS) properties, introducing edge weight normalization to calculate the connection strength between nodes. They defined the minimum connected dominating set efficiency (SEW) and excluded non-critical endpoints, further distinguishing node importance.
6 Machine Learning-based Ranking Methods
Machine learning-based ranking methods have transformed the identification of influential nodes in complex networks by leveraging advanced techniques such as information entropy, clustering coefficients, graph neural networks, and reinforcement learning to capture intricate structural patterns and relationships. These methods find applications in diverse domains, including social network analysis, information diffusion, network security, and infrastructure management as details in Table. 8. Compared to traditional approaches, machine learning enables more accurate and nuanced assessment of node importance, while also being capable of handling large-scale networks and capturing nonlinear dependencies that conventional methods may overlook. However, these approaches introduce new challenges, such as the requirement for extensive training data, dependence on model design and generalization capability, and limitations in computational resources.
6.1 Information Entropy
Entropy-based measures quantify the uncertainty or diversity in information propagation and have been widely applied to assess node centrality. Nikolaev et al. nikolaev2015efficient introduced entropy centrality, which evaluates a node’s potential for dissemination via discrete Markov processes. Specifically, a node’s entropy centrality is defined as
$H^{t}_{i}=-\sum_{j=1}^{N}(p_{ij}^{(t)}+p_{ij^{\prime}}^{(t)})\log_{2}(p_{ij}^{(t)}+p_{ij^{\prime}}^{(t)}).$ (21)
The term $p_{ij}^{(t)}+p_{ij^{\prime}}^{(t)}$ represents the probability that an object originates from node $v_{i}$ and that during $t$ time periods, the node itself is found to possess node $v_{j}$. Higher entropy indicates a more uniform spread, implying greater unpredictability in the information’s destination. Building on this foundation, Zareie et al. zareie2019influential incorporated Shannon entropy and Jensen–Shannon divergence to develop the Diversity Strength Ranking (DSR) and Extended DSR (EDSR) algorithms, which account for both neighbor dispersion and influence range.
Other studies have leveraged entropy in various contexts. Nitt et al. nitt2016using proposed a degree centrality mapping entropy (ME) to identify critical nodes through local interactions. Similarly, Fu et al. fu2015identifying defined global node information entropy based on the $k$-shell index. To address the issue in the VoteRank algorithm where all nodes are roughly assigned the same voting capacity and attenuation factor during voting, Guo et al. guo2020influential addressed limitations in VoteRank by introducing the EnRenew algorithm, which integrates connection-based entropy and an attenuation factor to update nodes’ propagation capabilities. In response to the susceptibility of road networks to cascading failures caused by traffic accidents or abnormal events, Xu et al. xu2018discovery applied entropy concepts in the Origin-Destination Entropy and Flow (ODEF) and Crossroad Ranking (CRRank) to evaluate node importance using vehicle trajectory data. Tulu et al. tulu2018identifying further developed a Community-based Mediator (CbM) method that combines internal and external density entropy with node degree to identify critical mediators in large-scale networks.
Network entropy is typically used to characterize the amount of information encoded in a network structure. Based on the assumption that removing more important nodes may lead to greater structural changes, Ai et al. ai2017node defined node importance as the change in network entropy before and after its removal. Wu et al. wu2023identify proposed $k$-order entropy combined with betweenness centrality to capture global and local information through different order neighbor dependencies. Based on K-shell, the internal and external effects of neighbor subnet structure are defined as information entropy and cross-entropy tong2023vital , respectively, to identify the information disseminator. Li et al. li2019key further extended this approach by incorporating mutual information to derive a centrality measure (EMI) that considers both direct and second-order neighbor effects, effectively quantifying redundant and complementary information flows. These entropy-based approaches offer a rigorous framework to evaluate node importance by measuring the uncertainty and diversity inherent.
6.2 Clustering Coefficient
The clustering coefficient is a crucial factor in determining the local influence of nodes. Chen et al. RN35 combined the number of node neighbors, the influence of neighbor nodes, and the clustering coefficient to propose the ClusterRank method for local node importance ranking, specifically defined as:
$s_{i}=c(v_{i})\sum_{j\in\Gamma(v_{i})}(k_{j}^{out}+1).$ (22)
The local clustering coefficient $c(v_{i})$ of node $v_{i}$ is defined as: $c(v_{i})=R(v_{i})/k(v_{i})(k(v_{i})-1)$, where $R(v_{i})$ represents the number of triangles formed by node $v_{i}$, and $\Gamma_{i}$ denotes the neighbors of node $v_{i}$. The term "+1" accounts for the degree contribution of the node itself to its neighboring nodes. When ranking nodes based on local centrality, traditional methods overlook the structural connections among neighbors. To address this, Gao et al. gao2014ranking proposed Local Structural Centrality (LSC), which integrates both the number of a node’s neighbors and their topological interconnectivity. By incorporating the local clustering coefficient, LSC quantifies the extent to which a node’s neighbors are interconnected, capturing structural nuances often ignored in purely degree-based metric. The local clustering coefficient reflects the density of connections among a node’s immediate neighbors, offering insights into its influence within the network. Nodes with similar local centrality values but higher clustering coefficients tend to exhibit stronger propagation capabilities, as their tightly connected neighborhoods facilitate more effective information diffusion. In addition, the traditional K-truss method, which decomposes networks based on edge clustering, may be affected by “false cores” when clustering characteristics are not accurately captured. Yang et al. yang2018identifying proposed an improved K-truss approach that incorporates both edge diffusivity and edge clustering coefficients, thus providing a balanced metric for identifying influential nodes. Complementarily, Zareie et al. zareie2020finding developed the ECRM method, which posits that nodes with lower correlation to their neighbors are more likely to be influential, as they can more effectively disseminate information to the network periphery.
To overcome the limitations of undirected graphs in representing conditional independence, Dablander et al. dablander2019node developed a causal framework based on directed acyclic graphs (DAGs). Assuming the Markov condition holds, they introduced the concept of "faithfulness" to ensure that the DAG accurately captures variable independencies. They further quantified causal impacts using Average Causal Effect (ACE) and Kullback-Leibler (KL) divergence. Their findings indicate that eigenvector centrality outperforms traditional metrics in predicting the causal influence of nodes within such networks. Liu et al. liu2022identifying11 integrated latent group information in power systems by developing the Group-Driven Framework for Identifying Critical Nodes (GDF-ICN) algorithm. This framework combines the structural and operational states of power systems with node group affiliations, iteratively optimizing group partitions around critical nodes. A fuzzy closeness metric is incorporated to enhance the description of group structure, while leveraging clustering effects to bolster the robustness of critical node identification. Wang et al. wang2022influential22 proposed an influential node evaluation algorithm grounded in information entropy theory. By defining metrics based on the number of triangles and edge weights, they introduced the concept of edge entropy weights—capturing differences in edge weights and their influence on neighboring nodes. This approach integrates k-core metrics to characterize local node influence, with higher entropy values indicating greater complexity and overall influence within the network. Zhang et al. zhang2013identifying presented a novel method focused on maximizing the range of influence. Their approach constructs an information transmission probability matrix, which quantifies the probability of information transfer between arbitrary node pairs.
6.3 Machine Learning
Machine learning offers a powerful framework for identifying influential nodes by reframing the problem as a regression task. These methods leverage supervised learning algorithms trained on node labels derived from epidemic model simulations and traditional centrality measures, capturing complex relationships between network structure and propagation potential. Zhao et al. zhao2020machine employed information infection vectors to encode topological characteristics and infection dynamics across diverse propagation scenarios. Using node labels obtained from SIR model simulations, they trained classifiers and regressors—including Naïve Bayes, decision trees, random forests, SVM, KNN, logistic regression, and MLP—to predict node importance. Yang et al. yang2021identifying36 introduced a network-aware local centrality index (NLC) that integrates topology with node embeddings. Applying DeepWalk to project network structures into a low-dimensional space, they incorporated third-order neighborhood information to refine influence estimation. Rezaei et al. rezaei2023machine356 further advanced this paradigm by engineering collective node features—spanning connectivity, degree, and coreness measures—and training an SVR model with an RBF kernel, demonstrating the efficacy of tailored feature representations in predicting node influence. However, existing models face limitations in feature representation and scalability. The development of deep learning and graph neural network approaches provides a powerful new framework for analyzing node attributes and quantifying influence.
Graph Conventional Network: Zhao et al. ZHAO202018 pioneered the deep learning framework InfGCN based on graph convolutional networks (GCN). InfGCN constructs node structural features—including degree, closeness, betweenness, and clustering coefficient—and builds a fixed neighbor network via breadth-first search (BFS). By training with actual infection rate vectors from SIR model simulations, InfGCN minimizes the error between predicted and actual infection rates to accurately estimate node influence. Similarly, Kumar et al. kumar2022influence23 adopted the Stru2vec node embedding algorithm ribeiro2017struc2vec along with a message-passing neural network (MPNN) to process node embeddings and SIR-based influence vectors, identifying the Top-$k$ influential nodes. Zhang et al. zhang2022new57 combined convolutional neural networks (CNN) with GNNs in their CGNN algorithm, which simplifies feature matrices and focuses on first- and second-order neighbors to label nodes via the SIR model and optimize a loss function for precise identification of critical nodes. Liu et al.  liu2022learning developed a ranking method based on self-supervised learning and graph convolutional models, leveraging both node features and network structural information.
Graph Embeddings: Wei et al. wei2024enimnr focused on reducing computational complexity. They selected candidate nodes based on shell connections and topological features, then used deep learning to generate low-dimensional vectors for calculating propagation dependency. Keikha et al. keikha2020influence proposed DeepIM, which extracts global and local structural features via CARE, constructs custom paths, and employs Word2vec to learn node feature vectors for measuring user relevance. Bouyer et al. bouyer2024maximizing presented ETIM, combining shell decomposition, graph embedding, and local structure features to decrease computation. Wu et al. YihangandHu2024 constructed node features using rule-equivalent similarity and Graph Convolutional Networks. Ahmad et al. ahmad2024learning proposed frameworks (LCNN) that merge CNNs with local node representations and multi-scale metrics. Rashid et al. rashid2024olapgn introduced the GN model, integrating deep learning and probabilistic properties via Graph Convolutional Networks to detect overlapping communities and key nodes. Xiong et al. xiong2024vital proposed the AGNN algorithm, fusing autoencoders with GNNs to generate topological feature embeddings through GCN and optimizing the ranking prediction model using listMLE. Yu et al. yu2020identifying developed methods RCNN that integrate GCNs and CNN-based adjacency features for effective critical node detection.
Graph Attention Network: Most existing node importance assessments in knowledge graphs fail to fully utilize available information and cannot capture node relationships and attribute features. To address this, Park et al. park2019estimating GENI uses predicate-aware attention mechanisms and flexible centrality adjustments to aggregate node and neighbor information across multiple GNN layers. It updates node embeddings via an aggregation function with a nonlinear transformation and maps these to importance scores using a scoring neural network. GENI also incorporates node in-degree centrality to adjust importance estimates. Define a GNN with $L$ layers of networks, where the $l$-th aggregates information from the l1l-1 th layer as follows,
$\overrightarrow{h}_{N(i)}^{l}\leftarrow\text{TransForm}^{l}\left(\text{Aggregate}\left(\left\{\left(\overrightarrow{h}_{j}^{l-1},w_{i,j}^{l}\right)|j\in N(i)\right\}\right)\right),$ (23)
$\overrightarrow{h}_{i}^{l}\leftarrow\text{Combine}\left(\overrightarrow{h}_{i}^{l-1},\overrightarrow{h}_{N(i)}^{l}\right),$
where $\overrightarrow{h}_{i}^{l-1}$ is the $v_{i}$ represents the information aggregation of the (l1)(l-1) th layer, each neighbor node $N(i)$ of node $v_{i}$ denotes the set of its neighbors (each initialized with a weight $w_{i,j}^{l}$). The aggregation function and nonlinear transformation function share parameters across all nodes. At each layer, node scores are updated through a weighted combination of their own value and neighbors’ values, with attention mechanisms dynamically adjusting connection influence. Initial node scores come from feature embeddings processed through a neural scoring network. GENI also incorporates a centralization score based on node in-degree, refined through scaling and offset adjustments. Finally, a nonlinear activation function integrates this score with learned representations. Furthermore, Park et al. park2020multiimport designed MultiImport, an end-to-end latent variable model to infer node importance from multiple sparse and possibly overlapping input signals. MultiImport uses edge-aware attention aggregation and node centralization adjustment within a GNN framework to learn robust node representations. It models the knowledge graph structure and uses clustering to group nodes with similar functional signals. Munikoti et al. munikoti2022scalable introduced ILGR, a scalable GNN-based framework that approximates node and link importance through attention-driven embedding. ILGR encodes local structure by aggregating neighborhood information, with an attention mechanism assigning adaptive weights based on structural roles. Ge et al. ge2024node proposed MFA-NIE, integrating structural, relational, and attribute features with an enhanced attention mechanism for feature extraction, and adjusts node importance using TOPSIS centrality. Chen et al. chen2024social developed a hierarchical graph-enhanced variational autoencoder that reconstructs neighborhood representations to learn social influence patterns, using hierarchical graph attention networks to capture influential neighborhoods and mitigate data sparsity. Liu et al. liu2022nie23 introduced NIE-GAT for inter-domain routing networks, mapping static node features to dynamic cascading failure impacts through graph attention networks. Additionally, Kou et al. kou2023identify23 refined GCN-based approaches by framing node influence prediction as a regression task and using a multi-head attention mechanism to integrate neighbor importance, with the IC model providing simulation-based ground truth for optimization.
Graph Contrast Learning: Recent advancements in contrastive learning have improved node importance estimation by integrating multi-view representations and attention mechanisms. Liu et al. liu2023node addressed the limitations of single-graph models by introducing the MCRL framework, which employs dual graph encoders and contrastive learning to derive multi-perspective node representations. This approach enhances the estimation of node importance by comparing embeddings generated through graph convolutional networks (GCN) and graph attention mechanisms (GAT). By incorporating a positive-negative sampling strategy, MCRL identifies nodes that maintain both high similarity and distinctiveness across multiple structural views—key indicators of importance. Building on this concept, Zhang et al. zhang2024label developed the Label-Contrastive Pretraining (LICAP) model, which employs hierarchical sampling to prioritize top-ranked nodes. LICAP utilizes a Predicate-Aware Graph Attention Network (PreGAT) to refine node embeddings, ensuring robust differentiation between influential and peripheral nodes. Meanwhile, Shu et al. shu2025node proposed AGCL, a contrastive learning framework for heterogeneous networks that leverages attention mechanisms to quantify both local and global node importance. By extracting structural features through contrastive learning and computing importance weights via attention mechanisms, AGCL offers a comprehensive approach to evaluating node influence in complex networks.
Graph Neural Networks: Huang et al. huang2022estimating introduced HIVEN, a GNN-based framework designed for heterogeneous information networks (HINs). By integrating a heterogeneous information aggregator and a meta-path-based mechanism, HIVEN effectively captures the structural and relational intricacies of HINs, improving both accuracy and efficiency in node ranking. Building on structural knowledge integration, Chen et al. chen2024deep developed SKES (Deep Structural Knowledge Exploitation and Synergy), which incorporates node centrality and similarity metrics to learn high-dimensional feature distributions. SKES refines importance predictions using optimal transport theory, enabling precise characterization of node significance within heterogeneous graphs. Similarly, Lin et al. lin2024node leveraged Large Language Models (LLMs) in the LENIE framework to enhance semantic representations in knowledge graphs. By employing a clustering-based triplet sampling strategy and node-specific adaptive prompts, LENIE enriches node embeddings, thereby improving the initialization of node importance estimation models. Higher-order network structures have also been incorporated into learning-based ranking methods. Zhao et al. zhao2023novel proposed HONNMA, a high-order neural network framework that employs motif attention to capture complex topological dependencies. This model encodes interactions through a weighted motif adjacency matrix and refines node embeddings using a skip-connected architecture, allowing for more robust node importance estimation. Addressing the Critical Node Detection Problem, Michos et al. michos2024critical applied Hopfield Neural Networks (HNN) to optimize network resilience. By reformulating as an energy minimization task, their approach identifies critical nodes by converging toward stable network states with minimal energy.
6.4 Reinforcement Learning
Reinforcement learning influences key node selection by using rewards and penalties to guide the process, treating the impact of nodes—whether on overall network performance or individual propagation capability—as the value incentive. In this framework, the objective is to identify a set of important nodes by either maximizing or minimizing a value-based objective function. To tackle high time complexity in critical node identification, Fan et al. fan2020finding proposed the FINDER framework. It minimizes the accumulated normalized connectivity (ANC) through a node deletion strategy. The ANC is defined as:
$R(v_{1},v_{2},\dots,v_{N})=\frac{1}{N}\sum_{k=1}^{N}\frac{\sigma(\mathscr{G}~\setminus\{v_{1},v_{2},\dots,v_{k})\}}{\sigma(\mathscr{G})}c_{v_{k}},$ (24)
where $c(v_{k})$ is the normalized removal cost of node $v_{k}$, and $\sigma$ measures network connectivity. This optimization process is modeled as a Markov Decision Process (MDP), defined by the following components: state space $\mathcal{S}$ represents the residual graph $\mathscr{G}^{\prime}$ after a sequence of node removals, action space $\mathcal{A}$ denotes the selection of a node $v\in V(\mathscr{G}^{\prime})$ for removal at each step, and reward function $r:\mathcal{S}\times\mathcal{A}$ is the reduction in ANC resulting from removing node $v$, incentivizing actions that lead to rapid network fragmentation. A strategy for selecting the next node to delete, updated iteratively via reinforcement learning algorithms as policy $\pi(a|s)$. The iterative process involves evaluating candidate actions based on their expected rewards, selecting the node that maximizes the reward signal, and updating model parameters using a greedy or approximate policy improvement strategy. To further reduce the computational complexity of heuristic algorithm-based node deletion, Tan et al. tan2024learning proposed a feature-sensitive graph attention network. This method represents nodes by their features and, combined with an adversarial double-depth network, constructs a flexible end-to-end framework for generalizing critical node identification across different scales.
To address information diffusion features in node importance analysis, several studies have integrated reinforcement learning with graph-based methods. Jaques et al. jaques2019social used a multi-agent reinforcement learning framework with counterfactual reasoning and measures like KL divergence to assess the impact of alternative actions on other agents, identifying influential nodes. Chen et al. chen2023identifying79 proposed DeepELE, combining graph embeddings with reinforcement learning in an SIS model. Each node is assigned a diffusion weight, and a state–action value function evaluates the network-wide infection impact after node removal, selecting key nodes by minimizing an objective function. Li et al. li2019disco introduced DISCO, combining network embedding with deep reinforcement learning for various node-related problems. Chen et al. chen2023touplegdd developed ToupleGDD, an end-to-end deep reinforcement learning framework coupling three graph neural networks with a double deep Q-network for parameter learning, addressing influence maximization while overcoming limitations like narrow IM formulations and low scalability. Ling et al. ling2023deep introduced DeepIM, generating latent representations of seed sets in a data-driven manner to capture diverse information diffusion patterns, and designing a novel objective function for optimal seed sets under flexible node centrality constraints. Li et al. li2022piano presented PIANO, merging deep reinforcement learning with network embedding to estimate node influence, providing a pre-trained model pool for direct application. Uthayasuriyan et al. uthayasuriyan2024adaptive introduced DERL, integrating differential evolution with DQN-based reinforcement learning, and Li et al. li2024adarisk proposed AdaRisk, a risk-adaptive deep reinforcement learning framework for detecting fragile nodes in uncertain graphs. Xu et al. xu2025influence developed HEDRL-IM, reformulating influence maximization in hypergraphs as a network weight optimization task solved via deep Q-networks, integrating evolutionary algorithms and propagation simulation. Zhu et al. zhu2025bigdn proposed BiGDN, employing bidirectional neighborhood aggregation and integrating deep reinforcement learning with multi-head attention for robust node representations. Ahmad et al. ahmad2025learning developed CoreQ, using K-core decomposition and Q-learning to identify candidate seed nodes.
Table 8: Comparison of Machine Learning-Based Ranking Methods.
Title Related Works Advantages / Disadvantages
Information Entropy nikolaev2015efficient ,  zareie2019influential ,  nitt2016using ,  fu2015identifying ,  guo2020influential ,  xu2018discovery ,  tulu2018identifying ,  ai2017node ,  wu2023identify ,  tong2023vital ,  li2019key ++ Robustness to various network structures. - Dependent on accurate data, hard for large or dynamic networks.
Clustering Coefficient RN35 ,  gao2014ranking ,  yang2018identifying ,  zareie2020finding ,  dablander2019node ,  liu2022identifying11 ,  wang2022influential22 ,  zhang2013identifying ++ Improves identification of influential nodes. - Potential oversimplification of network dynamics, reliance on local information.
Graph Conventional Network ZHAO202018 ,  kumar2022influence23 ,  zhang2022new57 ,  liu2022learning ++ Straightforward implementation, robustness in various network types. - Limited ability to capture complex patterns, potential inefficiency with large-scale networks.
Graph Embeddings wei2024enimnr ,  keikha2020influence ,  bouyer2024maximizing ,  YihangandHu2024 ,  ahmad2024learning ,  rashid2024olapgn ,  xiong2024vital ,  yu2020identifying ++ Improves accuracy for influential nodes. - Reliance on quality of embeddings.
Graph Attention Network park2019estimating ,  park2020multiimport ,  munikoti2022scalable  ge2024node ,  chen2024social ,  liu2022nie23 ,  kou2023identify23 ++ Heightened accuracy in ranking influential nodes, the ability to focus on important features. - Dependency on hyperparameter tuning, potential overfitting in small datasets.
Graph Contrast Learning liu2023node ,  zhang2024label ,  shu2025node ++ Effective identification of key nodes by leveraging differences between subgraph representations. - High computational cost, potential overfitting.
Graph Neural Networks huang2022estimating ,  chen2024deep ,  lin2024node ,  zhao2023novel ,  michos2024critical ++ Superior ability to capture complex patterns, adaptability to diverse network structures. - Dependence on extensive training data and difficulties in model interpretability.
Reinforcement Learning fan2020finding ,  tan2024learning ,  jaques2019social ,  chen2023identifying79 ,  li2019disco ,  chen2023touplegdd ,  ling2023deep ,  li2022piano ,  uthayasuriyan2024adaptive ,  li2024adarisk ,  xu2025influence  zhu2025bigdn ,  ahmad2025learning ++ Dynamic adaptation to network changes, improves decision-making over time. - Complexity in setting reward structures, reliance on extensive exploration and training periods.
7 Comprehensive Index Based Ranking Methods
Comprehensive index ranking methods evaluate node importance by integrating multiple network attributes and advanced analytical techniques. These methods combine various centrality measures, gravity models, and decision-making frameworks to overcome the limitations of single-indicator approaches and provide a more holistic assessment of node significance in complex networks. These indicators are widely applied across domains such as social network analysis, biological networks, transportation systems, and infrastructure management. The primary advantage of these methods is their ability to incorporate diverse structural and functional indicators, thereby offering a more nuanced characterization of node importance. However, they also introduce challenges, including increased computational complexity and careful selection and integration of relevant indicators.
7.1 Gravity Formula
The law of universal gravitation has been innovatively applied in network science. By defining the gravitational constant multiplied by the product of the masses of two nodes divided by the square of the distance between the nodes, the centrality and statistical indicators of nodes are considered as "mass", and the shortest distance between nodes is used as the distance, thus constructing a new framework for measuring node importance. Ma et al. RN14 combined the K-shell index of nodes and the shortest path between nodes to identify important nodes in the network through the gravitational calculation formula, defining the importance of node $v_{i}$ as:
$GM(v_{i})=\sum_{j\in\Psi}ks(i)\cdot ks(j)/d_{ij}^{2},$ (25)
where $ks(i)$ and $ks(j)$, denote the $k$-shell values of nodes $v_{i}$ and $v_{j}$, $d_{ij}$ is the shortest path between them, and $\Psi$ is the set of neighbors within a radius $r$ of node $v_{i}$. In parallel, Maji et al. maji2020influential addressed limitations stemming from empirically set free parameters in improved $k$-shell decomposition algorithms by considering the influence of different levels of neighbors, thereby enhancing the adaptability, efficiency, and accuracy of the model.
In addition, Li et al. RN16 proposed the Local Gravity Model (LGM), which extends the basic gravity model to include neighborhood information; nodes surrounded by neighbors with high $k$-shell indices are considered closer to the network’s center. Li et al. RN24 further modified the model by replacing the $k$-shell metric with the degree of neighboring nodes. Recognizing that a degree-based gravity model captures only local information. Liu et al. liu2020gmm introduced the eigenvector exponent of the adjacency matrix into the gravitational formula, proposing the Generalized Gravity Model (GMM) that incorporates both local and global network characteristics. Yang et al. RN13 utilized $k$-shell values to define an attraction coefficient between nodes, taking into account critical positional attributes of a node’s neighborhood. Li et al. RN20 measured local information using each node’s local clustering coefficient and clustering degree, thereby defining a node’s spreading ability as its mass derived from both neighbor information and node degree. In a related approach, Fei et al. fei2018identifying assessed node importance by defining mutual attraction based on the inverse-square law, though this method may overlook more complex inter-node relationships.
A single centrality measure often fails to fully capture the importance of nodes in a network. To address this, researchers have introduced weighted models based on gravitational principles, incorporating multiple attributes to assess node influence. Wang et al. RN28 enhanced the GM method by integrating $k$-shell value and degree centrality while further considering neighbor information to improve accuracy. Building on this approach, Li et al. RN26 proposed the MCCM method, which defines node importance as a weighted combination of degree, $k$-shell value, and eigenvector centrality. To ensure comparability across different networks, these metrics are normalized, with a balancing parameter introduced to prevent the $k$-shell value from disproportionately influencing smaller networks. Similarly, Yan et al. yan2020identifying adopted a multi-attribute decision-making approach, constructing a decision matrix based on multiple centrality measures and using entropy weighting to determine a node’s overall importance. Beyond centrality weighting, researchers have also addressed the issue of influence overlap among high-centrality nodes. Wang et al. wang2022identifying20 introduced the centripetal centrality index, which integrates node attributes and connectivity to reduce redundancy in influence propagation, albeit at a high computational cost. Yang et al. yang2023aigcrank17 developed the AIGCrank algorithm, combining gravitational centrality with a recursive ranking strategy that integrates the h-index, domain core centrality, and clustering coefficient, effectively distinguishing key nodes while mitigating influence overlap. Other approaches focus on probability-based propagation and structural optimization. Zhu et al. zhu2023identifying12 introduced the HVGC method, incorporating the h-index, node distance, and an exponential decay factor to accurately identify bridge nodes that connect different network regions, even if they exhibit low $k$-shell values. Additionally, Liu et al. liu2023entropy12 developed the SEGM and EGM models, respectively, leveraging information entropy to capture the diffusion capabilities of nodes more effectively.
To mitigate the high computational cost of shortest path calculations, researchers have proposed alternative approaches that approximate or replace shortest path metrics with more efficient methods. Zhao et al. zhao2022random21 employed random walks to estimate node distances, using degree centrality as a proxy for node influence. Extending this idea, Zhao et al. zhao2023estimating3 introduced the NEGM model, which replaces shortest path lengths with Euclidean distances derived from Node2Vec embeddings. Similarly, Shang et al. RN25 defined a probabilistic distance using Markov chain-based transition probabilities, capturing latent inter-node relationships without computing exact shortest paths. Beyond shortest path approximations, alternative gravity-based models have been proposed. Curado et al. curado2023novel introduced a method leveraging return-based random walks and effective distance to capture both static and dynamic network topologies. Yang et al. yang2023aogc refined interaction range definitions by incorporating an adaptive truncation radius and considering all-channel paths instead of shortest path constraints. Chen et al. chen2023identification proposed the DCGM model, integrating degree and average neighbor degree to enhance influence estimation. Further advancements address asymmetric interactions and multilayer structures. Meng et al. meng2025improved developed the Asymmetric Attraction Model (AAM), transforming adjacency matrices into asymmetric attraction matrices to capture directional dependencies. Xu et al. xu2024cagm introduced the CAGM model, incorporating influence probability, $k$-shell, and degree information for adaptive gravity-based estimation. Li et al. li2023improved123 extended gravity-based models to multilayer networks, proposing the APAMGM algorithm for multiplex transportation networks. Lü et al. lv2024improved introduced PRGC, integrating multi-PageRank centrality with weighted shortest path distances. Chi et al. chi2024measuring developed a semi-global centrality measure inspired by stellar models, incorporating variable propagation speeds and the gravitational slingshot effect to quantify node efficiency in information diffusion.
7.2 Multiple Metrics
To address challenges in identifying critical nodes arising from varied network structures and densities, researchers have developed composite centrality measures that integrate multiple metrics. For instance, Comin et al. comin2011identifying proposed a comprehensive centrality index that balances global and local characteristics by analyzing three different information propagation strategies. De Arruda et al. de2014role further explored dynamic propagation by proposing the Random Walk Accessibility (RWA) expansion indicator, based on nine centrality measures, to assess the ability of nodes to spread information in both spatial and non-spatial networks. To integrate multiple centrality measures, Hu et al. hu2015multi adopted linear discriminant analysis to integrate eigenvector, betweenness, closeness, degree centrality, and mutual information, developing a multi-metric evaluation algorithm. Hu et al. hu2015identifying then proposed the NICCM method, constructing a contribution matrix that defines importance by combining efficiency with weighted contributions.
Several studies have integrated multiple centrality measures to enhance prediction accuracy. Bucur et al. bucur2020top combined statistical classifiers into composite metrics, demonstrating high predictive power in SIR epidemic models. Wei et al. wei2022identifying24 developed a heterogeneous mean-field model incorporating betweenness, degree, H-index, and core degree, showing that betweenness-based immunization strategies are effective in BA networks, whereas degree-based strategies perform well in real networks. An et al. an2024novel introduced the DIRCI method, leveraging dynamic influence scope, network layer centrality, and community centrality for critical node identification in multilayer networks. Wu et al. wu2024hunting proposed radiation centrality, which models information dissipation using attenuation and scattering theories. Other approaches combine local and global structural features. Cao et al. cao2024dynamic introduced INLGC, integrating local network constraint coefficients with global community structure. Kopsidas et al. kopsidas2022identification merged centrality measures from different transport networks to assess metro station criticality. Wang et al. wang2024multi proposed a multi-factor information matrix centrality algorithm, incorporating node influence, neighbor influence, and feedback-based mutual influence. Lei et al. lei2024weighted developed the Weighted Information Index (WII), which utilizes second-order neighbor information to construct an information distribution vector. Ullah et al. ullah2024leveraging introduced NPIC, integrating neighborhood and path information, while Zhang et al. zhang2024critical employed the Comprehensive Voting Ranking (CVR) algorithm, incorporating mutual information and k-center clustering to identify critical nodes in urban rail networks. Several iterative and hybrid models have also been proposed. Lee et al. lee2024identifying developed the Balanced Iterative Influence (BII) algorithm, iteratively combining local structural information with global influence. Esfandiari et al. esfandiari2025collaborative introduced HNPR, which integrates $k$-shell and PageRank into a novel linear metric. Chen et al. chen2025sfimco proposed the SFIMCO method, distinguishing overlapping from non-overlapping nodes to quantify intra-community influence. Zhang et al. zhang2025novel introduced the SLCMF metric, integrating structural, social, and semantic factors while improving scalability via distributed subgraphs and shortest path approximations. Dempster-Shafer evidence theory has also been applied to node importance estimation. Mo et al. mo2015evidential developed Comprehensive Evidence Centrality (CEC), integrating degree, betweenness, and closeness centrality. Xu et al. xu2025lgp introduced the Local-Global-Position (LGP-DS) algorithm, which combines global, local, and positional attributes, leveraging information entropy to evaluate and aggregate their contributions to node importance.
To address the challenge of distinguishing nodes within the same $k$-shell layer, Sheikhahmadi et al. sheikhahmadi2017identification12 introduced the MCDE hybrid index, integrating node shell, degree, and entropy information to enhance accuracy and efficiency. Wang et al. RN51 ; wang2017ranking , Sheikhahmadi et al. sheikhahmadi2017identification , and Yang et al. yang2018ranking developed iterative removal-based indices that combine node position and factor indices to achieve efficient $O(n)$ complexity. Namtirtha et al. namtirtha2021best introduced the NGSC indicator, a weighted sum of $k$-shell values and neighbor degrees, adaptable to both dense and sparse networks. Hybrid approaches integrating local and global structure have also been explored. Ullah et al. ullah2021identification developed the Global Structure Model (GSM) to capture global node influence, while the Local and Global Centrality (LGC) metric ullah2021identifying12 combines $k$-shell entropy with neighbor degree sums. Hu et al. hu2022exploring209 proposed a neighbor similarity coefficient, integrating $k$-shell hierarchy and node degrees for improved influence evaluation. Wang et al. wang2022influential developed the ALSI method, incorporating both inherent and neighbor-driven influence. Mukhtar et al. mukhtar2023integrating introduced H-GSM, combining degree centrality and $k$-shell centrality to overcome single-metric limitations. Yang et al. yang2023identifying23 refined influence ranking using a structural iteration factor method, distinguishing nodes through iteration steps to mitigate propagation limitations. Qiu et al. qiu2024key proposed a hierarchical composite algorithm, integrating constraint coefficients from the salton index with improved $k$-shell for local-global ranking.
7.3 Topsis and Entropy-Weight
[…]
$A^{+}=\left\{v_{1}^{+},v_{2}^{+},\dots,v_{n}^{+}\right\}=\left\{\left(\max_{i}v_{ij}|j\in K_{b}\right)\left(\min_{i}v_{ij}|j\in K_{c}\right)\right\},$ (26)
$A^{-}=\left\{v_{1}^{-},v_{2}^{-},\dots,v_{n}^{-}\right\}=\left\{\left(\min_{i}v_{ij}|j\in K_{b}\right)\left(\max_{i}v_{ij}|j\in K_{c}\right)\right\},$
where $K_{b}$ and $K_{c}$ are benefit and cost sets. The Euclidean distances to these ideals, $S^{+}_{i}$ and $S^{-}_{i}$ were calculated, and the relative closeness $C_{i}=\frac{S^{-}_{i}}{S^{-}_{i}+S^{+}_{i}},i=1,\dots,m.$ A higher $C_{i}$ indicates that the node is closer to the positive ideal solution and farther from the negative ideal, thus signifying greater importance. Several studies have applied and extended this methodology. For instance, Liu et al. liu2015node used degree, betweenness, closeness, and $k$-shell values to construct and standardize a decision matrix for TOPSIS-based node evaluation. Hu et al. hu2016modified incorporated the SIR model to adjust attribute weights and optimize the TOPSIS model. Other researchers have tailored TOPSIS to specific network types. Li et al. li2023identifying23 combined resistance centrality with a weighted TOPSIS approach for weighted power grids. Dong et al. dong2022cpr34 defined a communication probability matrix based on shortest path length and number of paths between nodes, using relative entropy to measure differences in communication distributions in their TOPSIS system. These adaptations highlight TOPSIS’s flexibility and effectiveness in evaluating node importance across diverse network types.
The entropy weight method and empowerment method enhance the highest attribute in the comprehensive evaluation system, increasing the distinction of relative tightness. Chen et al. chen2021effects introduced an adjustable TOPSIS method to address the entropy weight method’s tendency to overemphasize attributes with high data diversity. This method incorporates a weight coefficient to modulate the influence of entropy weights on the evaluation outcomes. Ishfaq et al. ishfaq2022identifying45 employed the entropy weight technique to assign objective weights to each criterion and then applied TOPSIS to rank nodes. Yang et al. yang2018multi proposed a local centrality indicator that incorporates multi-layer neighbor and clustering coefficient information. When combined with grey relational analysis (GRA) and the SIR model, this indicator more accurately identifies key nodes compared to traditional weighted TOPSIS. Vega-Oliveros et al. vega2019multi introduced the multi-centrality index (MCI) for document keyword extraction by optimizing the combination of centrality indicators. Lu et al. lu2020node used relative entropy and gray relational degree to improve node evaluation in power networks. Zhang et al. zhang2022multi23 proposed a Multi-attribute Critic Network Decision Indicator (MCNDI) that integrates the H-index, closeness centrality, and other metrics using the objective CRITIC method for weighting. Ju et al. ju2021novel23 introduced the Multi-Criteria Compromise Ranking Method (VIKOR) combined multiple rail transit networks into a multi-layer regional network, determined centrality criterion weights using an objective method.
8 Information Propagation Based Ranking Methods
Ranking methods based on information propagation assess node importance by simulating and analyzing the spread of information or disease across the static network as illustrated in Table. 9. These methods offer valuable insights into dynamic spreading processes and are widely applied in social network analysis, public health disease control, and marketing strategies. Their main advantage lies in capturing both structural and dynamic aspects of diffusion, enabling a more realistic evaluation of node information maximization. However, they also face several challenges, including the difficulty of accurately modeling propagation dynamics, high computational demands for large-scale networks, and the need for domain-specific knowledge to interpret results effectively.
8.1 Diffusion Model
Diffusion-based ranking methods assess node influence by modeling information or disease spread. Macdonald et al. macdonald2012spreaders applied the SIR model to compare centrality metrics, demonstrating that eigenvector centrality effectively identifies high-impact spreaders above a critical infection threshold. Borgatti et al. borgatti2006identifying introduced the KPP-POS and KPP-NEG algorithms: KPP-POS optimizes information flow by maximizing average diffusion efficiency, while KPP-NEG disrupts network connectivity by removing key nodes, enhancing influence identification for targeted dissemination. Zhuge et al. zhuge2010topological proposed Topological Centrality (TC), iteratively updating node and edge weights to quantify influence, with a stable weight of 1 indicating centrality. Aral et al. aral2012identifying applied randomized experiments and a continuous-time proportional hazards model to distinguish individual influence from peer susceptibility. Addressing location-aware influence maximization, Li et al. li2014efficient developed approximation-ratio greedy algorithms: a boundary greedy method leveraging upper and lower bounds for Top-$k$ spreader identification, and a hint-based approach precomputing clues for efficient high-impact node detection in real-time dissemination.
To overcome the limitations of traditional centrality metrics in assessing the spreading capability of non-topologically dominant nodes, Lawyer et al. lawyer2015understanding introduced the Expected Force (EXP) metric, which incorporates entropy to model the distribution of infectiousness across multiple propagation rounds. Similarly, Chen et al. chen2015critical proposed three diffusion-based indicators—NEGD, diffusion speed, and diffusion scale—where NEGD, defined as the ratio of Expected Geodesic Distance (EGD) to Largest Geodesic Distance (LGD), quantifies overall propagation efficiency. Although these methods effectively identify influential nodes within local neighborhoods, they do not guarantee global optimality. Robinaugh et al. robinaugh2016identifying improved influence estimation by introducing single-step and multi-step expected impact measures for analyzing psychopathology networks. Bozorgi et al. bozorgi2016incim proposed the INCIM algorithm, which combines local propagation dynamics with global network structure to evaluate both inter- and intra-community influence under the linear threshold model. Meanwhile, Holme et al. holme2017three investigated outbreak size, vaccination effects, and information diffusion using SIR simulations. Collectively, these diffusion-based approaches simulate the spread of information or disease to assess node influence, integrating both local and global network characteristics to offer practical solutions for identifying critical nodes in diverse network environments.
Recent advances further enhance influence estimation. Yin et al. yin2023identifying47 leveraged Compressed Sensing Theory (CST) to frame node identification as sparse signal reconstruction, applying the SIR model for efficient target detection. He et al. he2019tifim proposed the TIFIM framework, combining descending iteration and top-advantage metrics for seed node selection. Tulu et al. tulu2019vital developed the Node Willingness and Influence (NWI) algorithm, integrating network structure and interaction frequency for key spreader detection. Zhong et al. zhong2021information introduced the Improved Information Entropy (IIE) method, leveraging higher-order neighborhoods and weighted entropy to refine node importance. Li et al. li2021estimating employed a Spearman-like correlation to enhance ranking precision, outperforming degree and betweenness centrality. Gong et al. gong2021probability proposed the Probability-Driven Structure-Aware (PDSA) algorithm, which updates influence estimates via graph traversal in the IC model. Wang et al. wang2023hgim introduced the HGIM method, constructing a heterogeneous propagation graph to integrate topology with diffusion cascades. To model dynamic influence, Mohammadi et al. mohammadi2024improved proposed the Two-Sided Sign-Aware Matching (TSM) framework, incorporating trust and reciprocity in signed networks. Wang et al. wang2024identifying introduced the Dynamic Propagation Probability (DPP) model, redefining neighbor influence within a three-hop neighborhood. Xu et al. xu2023novel developed the Local Propagation Probability (LPP) model, integrating hierarchical propagation influence across different orders.
For large-scale networks, Ai et al. ai2022identifying10 proposed the SPC method, which balances network connectivity with propagation precision by defining infection probabilities within a truncated radius. Zareie et al. zareie2020identification optimized key node selection using the Gray-golf algorithm, incorporating second-order neighbor effects for propagation efficiency. Chen et al. chen2020influential introduced the Dynamic Influence Seed Selection (DYISSE) method, integrating two-hop triangular influence measures for robust diffusion estimation. Fink et al. fink2023centrality introduced virus centrality, leveraging weighted directed networks to estimate propagation potential. Sun et al. sun2023finding further refined IDME by integrating neighborhood similarity and centrality. Ullah et al. ullah2023lss proposed the Local Structure System (LSS), which incorporates $k$-shell, degree, and triangle count to quantify influence. Finally, Corsin et al. corsin2025evidence developed an agent-based information diffusion model, demonstrating that source persistence has a greater impact on network consensus than source quantity, particularly in polarized environments.
8.2 Dynamical Sensitivity
Liu et al. liu2016locating has proposed the dynamic sensitive centrality (DSC) by integrating topological features and dynamic characteristics, considering a discrete time spreading model in which an infected node will infect its neighbors with a spreading rate $\beta$ and recover with a recovery rate $\mu$. Let $x(t)(t>0)$ denote the approximate value of the cumulative probability of nodes being activated between time step 1 and $t$, then x(t)x(t1)(t>1)x(t)-x(t-1)(t>1) is approximated as the probability of a node being infected at time step $t$. If $v_{i}$ is the unique initial infected node, then $x_{i}(0)=1$ and $x_{j\neq i}(0)=0$. In the first time step, $x(1)=\beta Ax(0)$, and for $t>1$: $x(t)-x(t-1)=\beta A[\beta A+(1-\mu)I]^{t-1}x(0)$, where $v_{i}$ is the identity matrix. Let $H=\beta A+(1-\mu)I$, then the cumulative probability of nodes being infected between time step 1 and $t$ can be approximated as
$x(t)=\sum_{r=2}^{t}[x(r)-x(r-1)]+x(1)=\sum_{r=0}^{t-1}~\beta AH^{r}x(0).$ (27)
Define $S_{i}(t)$ as the spreading influence of node $i$ at time step $t$. In the case where node $i$ is initially infected, it can be quantified as the sum of the infection probabilities of all nodes. The infection probability in the equation above is $x(t)=\sum_{r=0}^{t-1}~\beta AH^{r}e_{i}$ where $e_{i}=(0,\dots,0,1,\dots,0)^{T}$ is a vector with only the $i$-th entry being 1. Therefore, the spreading influence of node $v_{i}$ can be defined as $S_{i}(t)=[(\beta A+\beta AH+\dots+\beta AH^{t-1})^{T}L]$ where $L=(1,1,\dots,1)^{T}$ is a vector of size $n\times 1$. DS centrality directly measures node influence, capturing both structural and dynamic factors. Mo et al. mo2019identifying integrated multiple centrality metrics—including degree, betweenness, efficiency, and correlation—using the Dempster-Shafer (D-S) evidence theory. By treating these metrics as distinct probability assignments and applying evidence combination rules, they developed a comprehensive indicator for node importance. Du et al. du2020identifying introduced an Improved Topological Potential Model (ITPE) that leverages topological entropy to assign weights to centrality measures such as weighted degree, betweenness, closeness, and eigenvector centrality. By optimizing control parameters to minimize topological entropy.
Table 9: Comparison of Comprehensive Index-Based and Information Propagation Based Ranking Methods.
Title Related Works Advantages / Disadvantages
Gravity Formula RN14 ,  maji2020influential ,  RN16 ,  RN24 ,  liu2020gmm ,  RN13 ,  RN20 ,  fei2018identifying ,  RN28 ,  RN26 ,  yan2020identifying ,  wang2022identifying20 ,  yang2023aigcrank17 ,  zhu2023identifying12 ,  liu2023entropy12 ,  zhao2022random21 ,  zhao2023estimating3 ,  RN25 ,  curado2023novel ,  yang2023aogc ,  chen2023identification ,  meng2025improved ,  xu2024cagm ,  li2023improved123 ,  lv2024improved ,  chi2024measuring ++ Comprehensive consideration of multiple factors, flexibility in handling various network structures. - Potential complexity in parameter tuning, reliance on accurate index weightings.
Topsis and Entropy-Weight Multiple Metrics comin2011identifying ,  de2014role ,  hu2015multi ,  hu2015identifying ,  bucur2020top ,  wei2022identifying24 ,  an2024novel ,  wu2024hunting ,  cao2024dynamic ,  kopsidas2022identification ,  wang2024multi ,  lei2024weighted ,  ullah2024leveraging ,  zhang2024critical ,  lee2024identifying ,  esfandiari2025collaborative ,  chen2025sfimco ,  zhang2025novel ,  mo2015evidential ,  xu2025lgp ,  sheikhahmadi2017identification12 ,  wang2017ranking ,  sheikhahmadi2017identification ,  yang2018ranking ,  ullah2021identification ,  ullah2021identifying12 ,  hu2022exploring209 ,  wang2022influential ,  mukhtar2023integrating ,  yang2023identifying23 ,  qiu2024key ,  du2014new ,  liu2015node ,  hu2016modified ,  li2023identifying23 ,  dong2022cpr34 ,  chen2021effects ,  ishfaq2022identifying45 ,  yang2018multi ,  vega2019multi ,  lu2020node ,  zhang2022multi23 ,  ju2021novel23 ++ Holistic analysis by integrating multiple evaluation criteria, increases accuracy in identifying critical nodes, adaptability to different network types. - Difficulties in balancing and weighting diverse criteria effectively.
Information propagation method macdonald2012spreaders ,  borgatti2006identifying ,  zhuge2010topological ,  aral2012identifying ,  li2014efficient ,  lawyer2015understanding ,  chen2015critical ,  robinaugh2016identifying ,  bozorgi2016incim ,  holme2017three ,  yin2023identifying47 ,  he2019tifim ,  tulu2019vital ,  zhong2021information ,  li2021estimating ,  gong2021probability ,  wang2023hgim ,  mohammadi2024improved ,  wang2024identifying ,  xu2023novel ,  ai2022identifying10 ,  zareie2020identification ,  chen2020influential ,  fink2023centrality ,  sun2023finding ,  ullah2023lss ,  corsin2025evidence ,  liu2016locating ,  mo2019identifying ,  du2020identifying ++ Effective identification of nodes crucial for information flow, applicability to diverse network types. - Potential computational intensity, reliance on accurate modeling of propagation processes.
9 High-order Networks Based Ranking Methods
Unlike traditional pairwise connections as Graph in Figure. 3 (a) , high-order networks, such as simplicial complexes and hypergraphs as shown in Figure. 3 (b)(c). Centrality measures typically used in conventional networks can be extended by integrating higher-order relationships, leading to the development of novel node importance metrics. Recent advancements in hypergraph-based ranking methods have improved node importance estimation by integrating higher-order relationships. Kapoor et al. kapoor2013weighted extended degree centrality to hypergraphs, defining weighted degree centrality as the sum of weights of hyperedges connecting a node to others. Specifically, Given a hypergraph $H=(V,E)$, the weighted degree centrality of a node is defined as
$C^{h}_{d}(i)=\sum_{j=1,j\neq i}^{N}~\sum_{k=1,\{v_{i},v_{j}~\}\subset e_{k}}^{L}w_{k},$ (28)
where $w_{k}$ is the weight of the hyperedge $e_{k}$. Furthermore, Kapoor et al. kapoor2013weighted introduced a distinction between strong and weak ties based on Granovetter’s concept of tie strength, which enables more accurate node importance estimation. In a similar vein, Lee et al. lee2021betweenness introduced hyperedge betweenness centrality, adapting traditional betweenness measures for hypergraphs by converting them into bipartite graphs. This method refines the understanding of node influence in complex networks. Building on these methods, St-Onge et al. st2022influential developed an approximate master equation to study influence propagation in heterogeneous hypergraphs. They quantified node spreading potential using hyperedge cardinality and hyperdegree, identifying key seed nodes. Xie et al. XIE2023103161 proposed HADP, which prioritizes nodes with the highest initial fitness as seed set members. It constructs and updates an adaptive set of seed node neighbors, applying degree penalty terms across all nodes in large-scale networks. This approach applies degree penalty terms across all nodes in large-scale networks, thus avoiding the complexity of individually calculating the fitness of each node. Lee et al. Lee2023mmlmathXA and Mancastroppa et al. Mancastroppa2023HypercoresPL proposed decomposing hypergraphs into hypercores, where nodes belong to at least k hyperedges of size at least m. This method identifies localized diffusing nodes more effectively than core degree metrics.
Figure 3: Network Structure Illustration.
In terms of hypergraph eigenvector centrality, Benson et al. Benson_Austin_R extended eigenvector centrality to uniform hypergraphs using tensor Perron-Frobenius theory. Zhao et al. zhao2023ranking introduced higher-order centrality measures incorporating low-order network metrics. These metrics incorporate low-order network metrics to more accurately assess the relative importance of clusters in higher-order structures. Li et al. li2023influence proposed an electrostatic field-based algorithm to identify seed node sets in hypergraphs. Xie et al. xie2023vital introduced LHGC, which incorporates node degree and higher-order distance in a gravitational model. The lack of research on higher-order relationships in dynamic networks has been addressed by Zhao et al. zhao2023general developed centrality matrices for Pagerank, Hub, and Authority by adjusting weights of network motifs at different orders.
Beyond binary network models, simplicial complexes, composed of $k+1$ vertices forming $k$-dimensional simplexes such as edges, triangles, and tetrahedra, offer a natural framework to capture higher-order interactions in complex systems battiston2020networks . By incorporating simplexes of different orders, these structures allow a more detailed representation of the intricate relationships within networks, as highlighted by Goldberg goldberg2002combinatorial and Estrada et al. estrada2018centralities . Building on this framework, Serrano et al. serrano2020simplicial introduced generalized adjacency measures based on the intersection patterns among simplices, extending traditional centrality concepts to higher-order contexts. Their approach, which includes random walk-based measures over sequences of $k$-order simplexes, enables the characterization of structural features such as shortest path lengths, eccentricity, and network diameter granovetter1973strength . This methodology offers a comprehensive perspective for probing both global structure and local properties. Furthermore, several studies have adapted classical graph-theoretic methods to simplicial complexes. Notably, adjacency matrices defined over $k$-order simplexes have been employed to compute eigenvector centrality bonacich2007some and Katz centrality estrada2015first , thus providing tools to capture both local connectivity patterns and global topological features. Overall, simplicial complex models have emerged as a versatile and powerful framework for the analysis of complex networks, enabling deeper insights into their underlying interactions and dynamic behaviors.
10 Dynamic Networks Based Ranking methods
Recent research has focused on developing ranking methods for dynamic networks (Figure. 3(d)) to better capture the evolving nature of node interactions and improve node importance estimation. Kim et al. kim2012temporal introduced the time-ordered graph to simplify dynamic networks into static networks of directed flow. They expanded traditional centrality measures, such as degree, betweenness, and closeness to dynamic graphs using the following definitions:
$D_{i,j}(v)$ $=\sum_{t=i}^{j}2D_{t}(v),\quad C_{i,j}(v)=\sum_{i\leq t<j}~\sum_{u\in V\setminus v}\frac{1}{\Delta_{t,j}(v,u)},$ (29)
$B_{i,j}(v)$ $=\sum_{i\leq t<j}~\sum_{s\neq v\neq d\in V,\sigma_{i,j}(s,d)>0}\frac{\sigma_{t,j}(s,d,v)}{\sigma_{t,j}(s,d)},$
where $D_{i}(v)$ represents node $v$ at time $i$, $\Delta_{t,j}(v,u)$ is the shortest path length between nodes $v,u$ in the time interval $[t,j]$, $\sigma_{i,j}(s,d)$ refers to the shortest path length between nodes $s,d$ in the time interval $[i,j]$, and $\sigma_{t,j}(s,d,v)$ refers to the shortest path length containing $u$ in the path between nodes $s,d$ in the time interval $[i,j]$. Tsalouchidou et al. tsalouchidou2020temporal highlighted the significance of temporal betweenness centrality (TBC) in influencing information spread in dynamic networks. They proposed a bi-objective algorithm for the shortest-fastest path (SFP) in temporal graphs, combining spatial and temporal factors to define TBC. This metric is sensitive to spatial and temporal distances between vertices, enhancing the understanding of communication intermediaries in temporal networks. Similarly, Elmezain et al. elmezain2021temporal applied the Opshal algorithm to calculate degree centrality and shortest distances in weighted networks, defining metrics such as temporal degree and temporal closeness to assess node importance over time.
In the realm of multilayer temporal networks, Lü et al. lv2019eigenvector developed the ECMSim method, utilizing a fourth-order tensor to capture node similarity and PageRank for tensor-based networks. They further refined this approach with MTEIGBC and MTPRBC centralities lv2019pagerank , which employ cosine similarity to rank nodes across multiple dimensions including layers and timestamps. Expanding upon these concepts, Lü et al. lv2021hits proposed the MT-HITS centrality. This method incorporates inter-layer similarity coefficients and solves tensor equations to define centrality vectors across multiple dimensions. Bi et al. bi2021temporal adapted the gravitational model to dynamic networks through the Temporal Gravity Model (TGM). TGM defines temporal distance based on both network structure and temporal order, integrating static centrality measures with their dynamic extensions.
To address the limitations of fixed constants in multilayer coupled networks, Jiang et al. jiang2022identifying developed the enhanced similarity index (ESI) to assess interlayer coupling in multilayer networks, moving beyond fixed constants. By integrating node neighbors across two time layers and introducing an attenuation factor, they created a decay-based super adjacency matrix (ASAM) to identify key node pairs affecting the largest connected component in temporal networks. Taylor et al. taylor2017eigenvector integrated centrality matrices from different network layers into a joint $NT\times NT$ matrix, capturing node and time layer importance. They defined marginal and conditional centrality, leading to a time-averaged eigenvector centrality that emphasizes strong interlayer connections. Tao et al. tao2022sequential introduced sequential path trees, which require at least one time-related path between the root and any subnode. These trees incorporate propagation time, intermediate nodes, and reachable paths. A normalized feature matrix based on centrality in these trees quantifies node spreading ability. Chen et al. chen2024temporal extended the local clustering coefficient to dynamic networks, revealing hidden connectivity through the temporal local clustering coefficient. Pan et al. pan2011path highlighted the role of temporal distance in information dissemination speed in dynamic networks. Their findings highlight that dissemination speed is constrained by the average temporal distance between nodes, emphasizing the critical role of the temporal dimension in understanding network dynamics.
To track influential nodes in dynamic social networks, Song et al. song2016influential framed the problem as influence maximization in dynamic environments. They proposed an upper-limit exchange greedy algorithm, which optimizes influence coverage by replacing nodes in the previously identified set of influential nodes, thus enabling efficient continuous tracking as the network evolves. Building on these ideas, Huang et al. huang2017dynamic extended dynamic sensitivity to temporal networks by introducing Temporal Dynamic Sensitivity Centrality (TDC), which highlights that both the topology and dynamics of the network affect node spread. Additionally, Huang et al. huang2017centrality defined a super-evolution matrix to model temporal network structures, effectively reducing computational complexity by transforming the problem of calculating eigenvector centrality into the computation of eigenvectors from low-dimensional matrices. Additionally, Qu et al. qu2019temporal introduced the Temporal Information Gathering (TIG) process, which defines node importance based on four key factors: time information gathering depth, time distance matrix, initial information, and a weighting function. Furthermore, Arebi et al. arebi2022effective proposed a method based on temporal centrality measures to identify the minimum driver nodes set (MDS) that can fully control the network. Meanwhile, Li et al. Li10317077 proposed a temporal information fusion model, which successfully identifies critical nodes in dynamic power information networks from both a topological and informational perspective.
To address the high settlement cost of temporal Katz centrality, Zhang et al. zhang2024tatkc introduced the Two-stage Attention-based Temporal Knowledge Graph Convolutional Network (TATKC). This framework combines temporal attention mechanisms with MLP to predict node ranking, enhancing performance through normalization and neighbor node sampling strategies. Similarly, Yu et al. yu2020identifying22 constructed a feature matrix based on neighborhood information and used the SIR model to identify key nodes in temporal networks. By employing network embedding and machine learning, they proposed the MLI algorithm, which transforms the identification of key nodes into a regression problem. In addition, Yu et al. yu2023predicting proposed a framework that combines GCN and RNN to identify nodes with optimal propagation capabilities, considering the sequential topology of temporal networks. Furthermore, Wang et al. wang2024influential modeled node propagation as a hierarchical infection process, proposing Hierarchical Structure Influence (HSI), which quantifies a node’s propagation potential across the entire network.
11 Summary and Discussion
In this review, we have systematically examined the concepts and methodologies associated with identifying important, critical, influential, and key nodes in complex networks and systems. Existing research has been broadly categorized into three main classes: structurally important nodes, highly influential nodes in information spreading, and critical nodes under performance constraints. In static networks, traditional centrality measures, such as degree, betweenness, and clustering coefficient, have provided the foundation for assessing node importance. However, the advent of dynamic, multilayer, and higher-order network models has given rise to a new generation of metrics, including spatio-temporal coupling indicators, propagation-based dynamic centralities, and hypergraph-based approaches. These methods have deepened our understanding of node roles and spreading mechanisms, particularly in contexts where interactions evolve over time and across multiple layers. Furthermore, models that incorporate coupling effects, edge weight dynamics, temporal features, and information fusion have enabled more robust identification of critical nodes in complex, real-world settings. Such advances have demonstrated practical relevance across diverse domains, including social networks, biological networks, power grid networks, transportation networks, and geographical networks.
Despite these developments, significant challenges persist in accurately pinpointing key nodes within real-world networks characterized by heterogeneity and dynamic complexity. The following areas require fulture direction: (1) Dynamic Network Evolution Patterns: Existing methods often merely extend static indicators to dynamic networks, with insufficient study on the evolution patterns of key nodes and their spatial distribution in dynamic networks. This limits the exploration of potential patterns in complex systems. Additionally, researchers should focus more on the identifying and inferring higher-order structures and dynamics in higher-order networks. (2) Efficient Algorithm Development: As networks grow in scale and complexity, there is an urgent need for efficient algorithms with low computational overhead. Many existing heuristic methods for node removal lack scalability and are ill-suited to large, dynamically evolving systems. Research should focus on developing real-time tracking algorithms, influence estimation techniques, and control strategies tailored to dynamic and multilayer networks. (3) Integration of Machine Learning and Data-Driven Approaches: Recent advances in machine learning and data-driven methods, such as network embedding and deep learning, provide new perspectives for node importance identification. The emergence of novel network data structures like semantic text attribute graphs and knowledge graphs offers opportunities to combine traditional centrality measures with modern machine learning models. This integration can enhance prediction accuracy and adaptability in recommendation systems, databases, and other applications. (4) Cross-Domain Applications and Validation: Most current node importance measures are context-specific, limiting their generalizability. Developing a unified framework that accounts for structural, dynamic, and higher-order features is essential. Different domains (e.g., social networks, biological systems, power grids, transportation networks) have distinct definitions and requirements for key nodes. Future research should emphasize interdisciplinary collaboration, validate theoretical models with real-world datasets, and develop customized solutions for specific scenarios. In conclusion, the identification of key nodes in complex networks is a vibrant research area with significant theoretical and practical implications. Addressing the aforementioned challenges and future directions will contribute to a deeper understanding of complex systems and enable more effective solutions for real-world problems.
Table 10: Overview of article structure, method categorization, and main representative approaches.
Ranking Methods Categories Representative Methods
Neighbors-Based Degree Centrality DC RN3 , SLC RN4 ,LC ma2019quasi , ILC zhu2024identifying
K-shell Decomposition K-shell RN41 , MDD RN42 ,M-Centrality RN6
H-index H-index lu2016h , LH-index RN10 , EHC zareie2019ehc
K-truss K-truss 2008Trusses , Trust-core malliaros2016locating
Structural Hole SH yu2017critical , ISH yu2017identifying , VKC xu2018identifying , ICC liu2021identifying54 , SHKS zhao2023ranking21
Neighborhood Similarity HP wang2015link , ESCRM lu2022critical11 , DNS rao2022cbim , TC tong2024novel
Path Betweenness Centrality BC RN56 , ECC hage1995eccentricity , ASP RN59 , RASP RN61 ,LC song2015novel ,CNICC ventresca2015efficiently EOCI yang2020critical , , LASPN xiao2024new
Closeness Centrality $C$ freeman2002centrality ,EFF latora2001efficient ,GLR salavati2019ranking ,GLS sheng2020identifying
Random Walk RW RN7 , DRW iannelli2018influencers
Eigenvector Eigenvector Centrality EC bonacich1971factoring ,PCC ilyas2011identifying , LeadersRanK ahajjam2015leadersrank ,ECDS zhong2018identifying
Katz Centrality Katz katz1953new
Pagerank Algorithm Pagerank page1998pagerank ,AttriRank hsu2017unsupervised , TPR sheng2020identifying12 , EPR su2021identification67
Voting Mechanism LeaderRank RN33 , WLeaderRank li2014identifying ,Accumulated Nomination poulin2000dynamical ,VoteRank RN36
Control-Optimization Network Control R-PGME ding2017key , BCNs lu2019pinning ,  li2006controlling ,  yu2012distributed ,  yu2013synchronization ,  sun2025optimal ,  zhang2017effect ,  mo2019effects ,  jiang2023searching ,  cao2024synchronization ,  zhang2022necessary ,  gao2025effects
Maximization of Connected Component $C$-CNDP Boginski0007 , $C$-CNP arulselvan2011cardinality ,BPT li2011finding ,MOCNDP ventresca2015experimental ,DSFLA tang2020discrete
Minimization of Pair Connectivity ILP di2012branch ,EPC dinh2015assessing ,GRASP purevsuren2016heuristic , HHA addis2016hybrid ,MIQP chen2020critical ,HILPR shen2012discovery ,RGDC ventresca2014region , MSD yin2023mixed23 , BICND zhang2024evolutionary , EBC jiang2025identifying
Minimum Spanning Tree MST chen2004evaluation ,CNC hermelin2016parameterized , CPND addis2013identifying ,MCDS aringhieri2016local
Machine Learning Information Entropy DSR zareie2019influential , ME nitt2016using , EnRenew guo2020influential , ODEF xu2018discovery , CbM tulu2018identifying , EMI li2019key
Clustering Coefficient ClusterRank RN35 , LSC gao2014ranking , ECRM yang2018identifying , ACE dablander2019node , GDF-ICN liu2022identifying11
Machine Learning NLC yang2021identifying36 , SVR rezaei2023machine356 , InfGCN ZHAO202018 , CGNN zhang2022new57 , DeepIM keikha2020influence , ETIM bouyer2024maximizing , LCNN ahmad2024learning , AGNN xiong2024vital , RCNN yu2020identifying , NIE-GAT liu2022nie23 , MFA-NIE ge2024node , GENI park2019estimating , MultiImport park2020multiimport , ILGR munikoti2022scalable , MCRL liu2023node , AGCL shu2025node , MCRL liu2023node , LICAP zhang2024label , HIVEN huang2022estimating , SKES chen2024deep , LENIE lin2024node , HONNMA zhao2023novel
Reinforcement Learning FINDER fan2020finding , DeepELE jaques2019social , DISCO li2019disco , ToupleGDD chen2023touplegdd , DeepIM ling2023deep , PIANO li2022piano , DERL uthayasuriyan2024adaptive , AdaRisk li2024adarisk , HEDRL-IM xu2025influence , BiGDN zhu2025bigdn , CoreQ ahmad2025learning
Comprehensive Index-Based Gravity Formula GM RN14 , LGM RN16 GMM liu2020gmm , IGM RN28 , MCCM RN26 , AIGCrank yang2023aigcrank17 , CAGM xu2024cagm , HVGC zhu2023identifying12 , SEGM liu2023entropy12 , NEGM zhao2023estimating3 , DCGM chen2023identification , ACM meng2025improved , APAMGM li2023improved123
Multiple Metrics PRGC lv2024improved , IMC tan2006evaluation , KSC HuQingcheng2013 , RWA de2014role , NICCM hu2015identifying , DIRCI an2024novel , INLGC cao2024dynamic , CNT kopsidas2022identification , WII lei2024weighted , NPIC ullah2024leveraging , CVR zhang2024critical , BII lee2024identifying , HNPR esfandiari2025collaborative , SFIMCO chen2025sfimco , SLCMF zhang2025novel , CEC mo2015evidential , LGP-DS xu2025lgp , MCDE sheikhahmadi2017identification12 , NGSC namtirtha2021best
Topsis and Entropy-Weight TOPSIS du2014new , Entropy-TOPSIS dong2022cpr34 , GRA yang2018multi , MCI vega2019multi , MCNDI zhang2022multi23 , VIKOR ju2021novel23
Information Propagation-Based Diffusion Model KPP-POS borgatti2006identifying , TC zhuge2010topological , EXP lawyer2015understanding , NEGD chen2015critical , INCIM bozorgi2016incim , CST yin2023identifying47 , TIFIM he2019tifim , NWI tulu2019vital , IIE zhong2021information , PDSA gong2021probability , HGIM wang2023hgim , TSM mohammadi2024improved , DPP wang2024identifying , LPP xu2023novel , SPC ai2022identifying10 , DYISSE chen2020influential , IDME sun2023finding , LSS ullah2023lss
Dynamical Sensitivity DS liu2016locating , ITPE du2020identifying
High-order Networks DC kapoor2013weighted , BC lee2021betweenness , HADP XIE2023103161 , HyperCoreMancastroppa2023HypercoresPL , CEC Benson_Austin_R , HOC zhao2023ranking , EFS li2023influence , LHGC xie2023vital , Pagerank zhao2023general
Dynamic Networks TDC kim2012temporal , TBC tsalouchidou2020temporal , ECMSim lv2019eigenvector ; lv2019pagerank , MT-HITS lv2021hits , TGM bi2021temporal , ESI jiang2022identifying , D-Testsong2016influential , TDC huang2017dynamic , TIG qu2019temporal , MDSarebi2022effective , TATKC zhang2024tatkc , MLIyu2020identifying22 , OSAMyu2023predicting , HSIwang2024influential